### v1

# Point-MoE: Towards Cross-Domain Generalization in 3D Semantic Segmentation via Mixture-of-Experts

Abstract
While scaling laws have transformed natural language processing and computer vision, 3D point cloud understanding has yet to reach that stage. This can be attributed to both the comparatively smaller scale of 3D datasets, as well as the disparate sources of the data itself. Point clouds are captured by diverse sensors (e.g., depth cameras, LiDAR) across varied domains (e.g., indoor, outdoor), each introducing unique scanning patterns, sampling densities, and semantic biases. Such domain heterogeneity poses a major barrier towards training unified models at scale, especially under the realistic constraint that domain labels are typically inaccessible at inference time. In this work, we propose Point-MoE, a Mixture-of-Experts architecture designed to enable large-scale, cross-domain generalization in 3D perception. We show that standard point cloud backbones degrade significantly in performance when trained on mixed-domain data, whereas Point-MoE with a simple top-$k$ routing strategy can automatically specialize experts, even without access to domain labels. Our experiments demonstrate that Point-MoE not only outperforms strong multi-domain baselines but also generalizes better to unseen domains. This work highlights a scalable path forward for 3D understanding: letting the model discover structure in diverse 3D data, rather than imposing it via manual curation or domain supervision.
1 Introduction
In both language and vision, we have witnessed remarkable advances driven by two trends: the aggregation of massive heterogeneous datasets [59, 64, 5, 57, 9, 89], and the deployment of models of ever-increasing capacity trained on them in a unified manner [8, 18, 55, 39, 25, 49]. These models succeed not because they were crafted for a specific domain, but because they were trained to learn underlying patterns across a wide variety of datasets. However, 3D point cloud understanding has yet to follow the same trajectory of progress, despite its wide applications in robotics[13, 50, 84], autonomous systems[15, 14, 44], augmented reality[90, 75], and embodied intelligence[77, 71]. Taking the 3D semantic segmentation task as an example, although diverse 3D datasets exist—such as ScanNet [21], SemanticKITTI [6], nuScenes [11], and ASE [3]—each covers only a narrow subset of the real-world diversity in 3D assets. Point clouds are generated from a variety of pipelines: RGB-D sensors [12], terrestrial and airborne LiDAR [33, 51], and multi-view stereo [83]. Each source brings its own characteristics—differing point densities, coverage patterns, reconstruction artifacts, and semantic biases (Fig. 1 left). The result is a 3D ecosystem of rich, yet siloed data sources, where models trained in one silo often fail to generalize across others.
A natural response to address this challenge is to train across such silos. For example, one might naively train the state-of-the-art Point Transformer V3 (PTv3) [78] (Fig. 1a) on mixed-domain 3D datasets. However, such models struggle to reconcile heterogeneous characteristics of the data, even with increased network capacity (Tab. 1). To mitigate this, recent methods introduce domain-aware components. Point Prompt Training (PPT) [80] (Fig. 1b) modifies PTv3 by assigning domain-specific normalization layers. Similarly, One-for-All [74] achieves improved 3D detection performance by employing a lightweight classifier to infer dataset labels, then applying domain-specific adapters. While these methods show notable gains, they rely on access to domain labels during training and exhibit performance degradation on unseen domains. A likely cause is that adapting only the normalization parameters may be insufficient to fully capture the complex variations required for scalable, cross-domain generalization.
In this work, we hypothesize that Mixture-of-Experts (MoE) architectures [35, 38] are well-suited for cross-domain joint training. We introduce Point-MoE, a sparse MoE model built upon PTv3 [78]. As illustrated in Fig. 1c, Point-MoE replaces each feed-forward layer in PTv3 with a MoE layer comprising of multiple expert networks (implemented as feed-forward MLPs) and a router that dynamically selects a sparse subset of experts for each input token. Our design is motivated by three key insights. First, MoE provides a natural inductive bias for handling diverse inputs by encouraging expert specialization and learning token-to-expert routing [17, 82]. This allows a single model to be trained on mixed-domain data without requiring access to domain labels. Second, compared to normalization-based domain adaptation methods like PPT [80] and One-For-All [74], MoE offers a significantly larger and more expressive representation space, enabling better modeling of domain-specific variations. Third, sparse MoE architectures improve computational efficiency during both training and inference, and have emerged as a standard design for scaling large models [41, 28, 26, 58, 22]. We extend these empirical benefits of MoE, previously established in language and vision, to the domain of 3D point cloud understanding, with a particular focus on 3D semantic segmentation.
Figure 1: Overview of cross-domain training architectures. Point clouds exhibit diverse characteristics across domains. (a) Naively training Point Transformer V3 (PTv3) [78] on multi-domain data leads to degraded performance within each domain. (b) Point Prompt Training (PPT) [80] addresses this by adding domain-aware normalization parameters. (c) Our proposed Point-MoE tackles this challenge with Mixture-of-Experts (MoE), enabling dynamic expert specialization across domains. Domain labels can optionally be provided to the router to guide the expert selection.
We conduct extensive experiments across a diverse mixture of 3D datasets, spanning indoor and outdoor scenes, synthetic and real data, and multiple sensor types. To comprehensively assess the effectiveness of Point-MoE, we evaluate it under both domain-agnostic (i.e., without access to domain labels) and domain-aware training regimes. Our results show that Point-MoE consistently adapts well to both settings, achieving strong generalization to unseen domains while outperforming competitive baselines, with improved training (Fig. 2) and inference efficiency (Tab. 1 and 2). We also conduct a thorough ablation study analyzing key design choices, including the number of experts, sparsity level, and the placement of MoE layers (Tab. 3). Remarkably, even without domain supervision, Point-MoE learns to route input data meaningfully and to specialize experts organically—emerging from the data distribution itself (Fig. 3 and 4).
To the best of our knowledge, this work presents the first systematic exploration of the Mixture-of-Experts architecture for 3D point cloud understanding in the context of cross-domain generalization. These findings point toward a scalable path forward for 3D perception: instead of designing separate models for each dataset or domain, we should aim to develop unified systems that can adapt across the full spectrum of 3D data sources. In this regard, Point-MoE emerges as a promising and flexible framework for bridging domain gaps in large-scale 3D semantic segmentation.
2 Related Work
Point Cloud understanding.
Deep learning backbones for 3D scene understanding can be broadly categorized by how they process point clouds: projection-based methods [15, 40, 42, 62], voxel-based methods [48, 61, 19, 29], and point-based methods [52, 53, 67, 86]. More recently, transformer-based architectures operating on point clouds have emerged as strong contenders in this space [31, 79, 87, 78]. Among them, Point Transformer V3 (PTv3) [78] represents the current state of the art, introducing a novel space-filling curve based serialization scheme, that transforms the unstructured 3D points into a 1D sequence, enabling efficient transformer-based modeling. A recent series of works have explored scaling such point cloud models to the multi-dataset setting using domain-aware training strategies. Point Prompt Training (PPT)  [80] demonstrates that using domain-specific normalization layers during pretraining improves 3D segmentation performance across multiple datasets. In 3D detection, One-for-All [74] adopts a sparse convolutional network with domain-specific partitions and normalization layers to boost multi-domain performance. Rather than relying on such domain-specific architectural branches [80, 74], we propose a unified sparse expert framework that learns to specialize dynamically across various domains, without requiring explicit dataset labels during training or inference.
Mixture-of-Experts.
Mixture-of-Experts (MoE) [35, 38] aims to increase model capacity without a proportionate increase in compute costs by levereaging conditional computation – large portions of the model are only activated conditioned on certain input samples [60]. In NLP, MoE has enabled scaling up models from hundreds of millions to hundreds of billions (and beyond) parameters: GShard scales to 600 B parameters [41], Switch Transformer to 1.6 T parameters [28], ST-MoE to 260 B parameters [91], and GLaM to 1.2 T parameters [26], all at roughly the same training cost as a dense model one-tenth the size. V-MoE [58] applies this recipe to vision, demonstrating competitive accuracy at 15 B parameters with batch-priority routing. More recently, Mixtral 8x7B [37] and DeepSeek models [22, 23, 24] all employ Mixture-of-Experts for large language model pretraining.
Apart from being a powerful tool for scaling up model capacity, MoE is well-suited to training on data drawn from heterogeneous domains, because the experts naturally tend to self-organize on a per-domain basis. Early work shows gains in multi-source sentiment and POS tagging [30], domain-adaptive language modeling [32], automatic chart understanding [82], and multilingual translation [17]. Mod-Squad [16] introduces a modular mixture-of-experts framework for multi-task learning, optimizing task-expert matching to balance cooperation and specialization across tasks. DAMEX [36] applies this to visual domains, while MoE-LLaVA [45], DeepSeek-VL [47], and DeepSeek-VL 2 [81] report similar benefits in vision-language models. To our knowledge, ours is the first attempt at extending MoE to address multi-dataset training and cross-domain generalization for 3D point clouds.
3 Point-MoE for Cross-Domain Point Cloud Understanding
Task definition. Given a 3D scene represented as a point cloud $\mathcal{P}=\{p_{i}\}_{i=1}^{n}$, where each point $p_{i}\in\mathbb{R}^{3}$ denotes coordinates, semantic segmentation aims to predict a single class label $\hat{y}_{i}\in\mathcal{C}$ to each point $p_{i}$ from a fixed label set $\mathcal{C}=\{c_{1},c_{2},\ldots,c_{m}\}$. Multi-domain data are utilized during training. Let the domain containing $N$ point clouds be $\mathcal{D}=\mathopen{}\mathclose{{}\left\{\mathopen{}\mathclose{{}\left(p_{i},y_{i}}\right)_{i=1}^{n}}\right\}^{N}$. Then the complete training set across $d$ domains can be represented as: $\mathopen{}\mathclose{{}\left\{\mathcal{D}_{k}}\right\}_{k=1}^{d}$. The goal of multi-domain joint training is to learn a unified model over all domains that minimizes prediction error across $\mathopen{}\mathclose{{}\left\{\mathcal{D}_{k}}\right\}$, while also generalizing effectively to domains unseen during training.
We adopt a minimal Mixture-of-Experts design that closely follows the standard architecture prevalent in the recent NLP literature. This simplicity allows us to leverage existing scalable Transformer-based MoE implementations with minimal modification. For the base model, we use Point Transformer V3 (PTv3) [78], a state-of-the-art architecture for point cloud understanding. Below we introduce the architecture of Point-MoE in detail.
Mixture-of-Experts layer. The core of MoE is to route input tokens dynamically to specialized subnetworks, referred to as experts, via a lightweight gating mechanism. Given an input feature vector $\mathbf{x}\in\mathbb{R}^{d}$, each MoE layer contains $N$ experts $\{f_{i}\}_{i=1}^{N}$ and a gating network $G:\mathbb{R}^{d}\rightarrow\mathbb{R}^{N}$. For each input $\mathbf{x}$, the MoE layer output is computed as a weighted sum over the top-$k$ activated experts:
$\mathrm{MoE}(\mathbf{x})=\sum_{i\in\mathcal{S}_{\mathbf{x}}}G_{i}(\mathbf{x})f_{i}(\mathbf{x})$ (1)
where $\mathcal{S}_{\mathbf{x}}$ denotes the set of top-$k$ experts selected for input $\mathbf{x}$, and $G_{i}(\mathbf{x})$ represents the normalized gating weight assigned to expert $i$. In practice, the gating function is typically implemented via a linear projection followed by a sparse softmax, ensuring activation of only a small subset of experts, with $k\ll N$. Although auxiliary losses for balancing expert load are common in prior work [60], we empirically observed them to be unnecessary and therefore excluded them from our final setup.
Integration into PTv3. To tailor the MoE architecture specifically for 3D point cloud data, we integrate the MoE mechanism directly into PTv3 by replacing the shared linear projection layers in PTv3’s attention module with an MoE layer. The original linear projections mapping input features into key, query, and value vectors within the attention mechanism are replaced by MoE layers, while all subsequent computations remain unchanged. This integration allows the MoE experts to specialize naturally in domain-specific transformations of point cloud data. Fig. 1c provides a visual overview of the modified architecture. It is crucial to insert the MoE layer before the post-normalization; otherwise, we observe a significant reduction in performance gains.
Language-guided classification. To bridge label discrepancies across datasets, we project per-point features to a shared language space using CLIP [56] text embeddings, following PPT [80]. This enables supervision via class names, without relying on dataset labels. For instance, the class “pillow” is missing in ScanNet (grouped as “other”) but is explicitly labeled in Structured3D. Our model aligns point features with the CLIP [56] embedding of “pillow,” allowing knowledge to transfer across domains even in the absence of direct supervision.
Domain-aware gating. Our preliminary experiments show that while basic Point-MoE naturally specializes experts implicitly, providing explicit domain information could further guide and accelerate expert specialization. To explore this, we introduce a dataset-aware gating variant termed Domain-Aware Point-MoE. In this variant, we explicitly associate each training domain $\mathcal{D}_{i}$ with a unique, learnable domain embedding vector $d_{i}\in\mathbb{R}^{d}$. Before routed through the gating mechanism, each input feature $x_{t}$ is concatenated with the corresponding domain embedding: $G_{t}(d)=\mathrm{softmax}(W\cdot(x_{t}\oplus d_{i}))$, where $W$ denotes a learnable linear projection, and the concatenation operation explicitly conditions gating decisions on the domain identity. This explicit conditioning encourages faster and more clearly delineated expert specialization aligned with dataset-specific characteristics observed during training.
Domain randomization for robust gating. While domain-aware gating allows experts to specialize based on known domain identities, this approach assumes access to explicit domain labels at inference time. In practical scenarios, such labels are often unavailable—especially when deploying models in the wild, where domain boundaries may be ambiguous or entirely unknown. While a domain detector can be employed to approximate the domain label by mapping unseen domains to the closest known ones, such detectors may introduce errors or fail to capture nuanced domain shifts. This motivates the need for a more robust mechanism that does not rely on explicit domain identification at test time. To address this, we introduce a domain randomization strategy that enables the model to generalize to unseen or mixed-domain inputs without performance degradation on seen domains.
Specifically, we define a learnable generic domain embedding and randomly assign it to 20% of the training points, replacing their true domain-specific embeddings. This forces the gating mechanism to learn domain-agnostic routing behavior, even in the presence of domain ambiguity. The remaining 80% of points retain their original domain embeddings, ensuring that the model still benefits from explicit supervision where available. In line with prior work [70, 17], we find that domain randomization improves generalization to unseen domains while preserving strong performance on those seen during training.
Training objective. We use Mixture-of-Experts (MoE) in a supervised cross-domain setting, where the goal is to train a unified model $f(x;\theta)$ across $n$ domains $\{\mathcal{D}_{i}\}_{i=1}^{n}$, each containing samples $(X,Y)\sim\mathcal{D}_{i}$. The objective is to minimize the average loss across domains:
$\min_{\theta}\sum_{i=1}^{n}\mathbb{E}_{(X,Y)\sim\mathcal{D}_{i}}\;\mathcal{L}(f(X;\theta),Y)$ (2)
We consider two batching strategies during training: (1) Single-domain batching, where each batch is sampled from only one domain $\mathcal{D}_{i}$ at a time. (2) Mixed-domain batching, where each batch contains samples from multiple domains. Our implementation is based on the official open-source code of PTv3 released by Pointcept [20].
4 Experiments
We conduct comprehensive experiments to evaluate Point-MoE for cross-domain 3D semantic segmentation. First, we describe baseline methods and experimental setups (Sec. 4.1). Then, we present detailed empirical results under two primary joint training scenarios: (1) indoor-only, and (2) combined indoor-outdoor training (Sec. 4.2). Finally, we provide ablation studies that highlight key design choices of the MoE architecture (Sec. 4.3) and detailed analysis of MoE behaviors (Sec. 4.4).
4.1 Experimental Setup
Baseline methods. Since our proposed Point-MoE framework is built upon PTv3 [78], we adopt PTv3 and PPT [80] as primary baselines. PPT, originally designed for dataset-specific BatchNorm  [34] and LayerNorm [4] layers and fine-tuning, is adapted here for joint training. Because PPT requires dataset labels at inference time, we follow One-for-All [74] to train a lightweight domain classifier (achieving over 95% accuracy), which we also use for Point-MoE under the domain-aware setting.
To ensure a fair comparison, we begin by benchmarking the dense PTv3 model under the domain-agnostic joint training setting, systematically varying key design factors such as mixed-batch composition and normalization strategy. In this setup, models are trained across multiple domains without access to dataset labels, and no label information is provided at inference time. This setting reflects a realistic deployment scenario where domain metadata is unavailable. For the domain-aware joint training setting, we compare to the PPT + One-for-All baseline. Here, models are trained with access to dataset labels. During inference, we either provide predicted labels using a trained domain classifier or assign a “generic” class to simulate deployment in unknown or mixed domains. This allows us to isolate the value of domain conditioning and assess the robustness of Point-MoE under both domain-aware and domain-agnostic regimes.
Datasets. Experiments are conducted on three indoor datasets—ScanNet [21], S3DIS [2], and Structured3D [88]—and evaluated on both seen and unseen domains, including Matterport3D [12]. We then evaluate the best-performing Point-MoE and baseline configurations under a more challenging indoor-outdoor setting, incorporating outdoor datasets from nuScenes [11], SemanticKITTI [7], and Waymo [65]. Voxel sizes are standardized to 0.02m for indoor scenes and 0.05m for outdoor scenes. Throughout, we denote model capacity using -S (small) and -L (large) suffixes for all architectures, including PTv3, Point-MoE, and PPT.
4.2 Experimental Results
Table 1: Performance of Joint Training for Indoor 3D Semantic Segmentation. All models are jointly trained on three indoor datasets and evaluated under two settings. In Domain-agnostic joint training, models are evaluated using only the point cloud and the corresponding dataset-specific label set. In Domain-aware joint training, a predicted scene label from a domain classifier is provided at inference time to guide the model. “-G” indicates that model utilizes domain randomization and contains a learnable “generic” domain label during training. All results are reported as mean Intersection-over-Union (mIoU).
Methods Params Activated Mix Layer. ScanNet Structured3D S3DIS Matterport3D
Seen Unseen
Val Val Test Val Val Test
Single-dataset training
PTv3-ScanNet [78] 46M 46M 75.0 19.8 20.9 8.8 38.9 44.5
PTv3-Structured3D [78] 46M 46M 21.2 81.5 79.4 7.1 21.0 24.0
PTv3-S3DIS [78] 46M 46M 4.9 3.5 5.3 67.6 5.0 3.4
PTv3-Matterport3D [78] 46M 46M 53.1 22.9 23.7 9.8 54.1 56.0
Domain-agnostic joint training
[…]
Domain-aware joint training
[…]
Table 2: Cross-domain 3D Semantic Segmentation across Indoor and Outdoor Scenes. We evaluate the best-performing models from the indoor-only setting in a more challenging joint training setup that spans both indoor and outdoor datasets. This setup better reflects real-world deployment scenarios, requiring models to generalize across diverse domains.
ScanNet Structured3D S3DIS nuScenes SemKITTI Matterport3D Waymo
Indoor Outdoor Indoor Outdoor
Seen Unseen
Methods Params Activated Mix Layer. Val Val Test Val Val Val Val Test Val Test
Domain-agnostic joint training
[…]
Domain-aware joint training
[…]
Single-dataset training. We begin by evaluating the standard setting in which models are trained and evaluated on individual datasets. Note we train end evaluate PTv3 with language-guided classification. As shown in Tab. 1, PTv3 achieves strong in-domain performance, but its performance drops significantly on unseen domains. For instance, a PTv3 model trained on S3DIS achieves only 3.5% on Structured3D and 4.9% on ScanNet. While domain-specific models are not expected to generalize across domains or categories, these results underscore the brittleness of single-dataset training. This motivates the need for cross-domain training strategies that expose the model to diverse data and promote broader generalization.
Domain-agnostic joint training. In the indoor-only setting, Point-MoE outperforms all variants of PTv3 in most domains and performs competitively on S3DIS. Referring to Tab. 1, Point-MoE-S achieves 75.4% on ScanNet and 69.7% on S3DIS, outperforming PTv3-S by +4.1% and +13.8%, respectively. Remarkably, Point-MoE-S activates only 52M parameters compared to 97M in PTv3-L, yet still surpasses PTv3-L by +0.8% on ScanNet and +8.4% on S3DIS. In addition, Point-MoE-L reaches 75.8% on ScanNet and 73.0% on Structured3D, significantly ahead of PTv3-L. On the unseen Matterport3D domain, Point-MoE-L obtains 41.6% on validation set and 47.8% on test set, outperforming PTv3-L by +1.4% and +1.2% respectively while activating only 59M parameters at inference. In the more challenging indoor and outdoor setting, Point-MoE-L again leads across the board with a few being competitve. Referring Tab. 2, Point-MoE-L achieves 69.2% on S3DIS validation and 70.2% on test split, outperforming the best dense baseline by +5.4% and +6.3%, respectively. On outdoor datasets, Point-MoE-L achieves 72.9% on nuScenes exceeding the dense baseline by +2.6%. Notably, on the unseen datasets, Point-MoE-L outperforming all dense PTv3 variants by a clear margin. These gains are achieved while activating nearly half the parameters of PTv3-L, underscoring the efficiency and scalability of sparse expert routing.
Domain-aware joint training. We further evaluate Point-MoE in the domain-aware setting, where models are trained with access to dataset labels. In the indoor-only setting, both domain-aware Point-MoE and its generic variant (Point-MoE-G) surpass all strong baselines: variants of PPT [80] + One-for-All [74]. As shown in Tab. 1, Point-MoE-G-L achieves 76.7% on ScanNet, 68.0% on Structured3D, and 67.0% on S3DIS, outperforming the best PPT + One-for-All baseline by +2.0%, +3.1%, and +2.3%, respectively. On the unseen Matterport3D dataset, it reaches 41.8% on the validation split and 47.8% on the test set, again outperforming all baselines. Notably, using only 20% of training data tagged as “generic” does not degrade performance on seen domains.
In the indoor-and-outdoor setting, Point-MoE-G-L continues to outperform all baselines across the board. Referring to Tab. 2, it achieves 77.0% on ScanNet, 72.1% and 74.2% on Structured3D (val/test), 62.4% on S3DIS and 72.7% on nuScenes—surpassing the best baseline by +0.3%, +4.6%, +4.1%, +1.0% and +0.5%, respectively. On the unseen domains, including Matterport3D and Waymo, Point-MoE-G-L delivers the strongest generalization. Crucially, it performs on par with domain-aware Point-MoE on seen datasets while achieving superior accuracy on unseen ones. These results validate that domain randomization with a generic label can enhance robustness without compromising in-domain performance. We still observe a performance gap between cross-domain and single-dataset training in cases like Structured3D [88]. Single-dataset models can adapt to domain-specific cues and distributions, while Point-MoE must balance capacity across domains. As a result, its performance may trail on large, homogeneous datasets, but remains competitive overall due to its generalization ability.
4.3 MoE Design Ablation
Table 3: Ablation of Point-MoE Design and Hyperparameters. We conduct a systematic ablation on key design and hyperparameter choices of Point-MoE across four indoor multi-domain datasets: ScanNet (Scan), Structured3D (S3D), S3DIS, and Matterport3D (Mat).
[…]
$10^{-4}$ 72.3 64.1 67.9 39.5
$10^{-3}$ 71.6 58.8 67.4 37.4
(a) Auxiliary load balancing loss. Removing the auxiliary loss consistently improves performance.
[…]
top 1 74.4 64.8 65.5 39.0
[…]
top 3 73.8 62.4 67.8 39.7
(b) Top-$k$ expert selection. Best performance is achieved with $k{=}2$ experts.
[…]
LN. 70.8 56.9 67.3 37.2
RMS. 45.3 36.5 51.1 24.6
(c) Normalization. BatchNorm yields strongest performance; RMSNorm degrades results.
[…]
dec. 74.0 64.9 66.6 38.1
[…]
(d) MoE position. Inserting MoE in all layers yields the best results.
[…]
(e) Activation functions. ReLU performs the best.
[…]
(f) Expert width. H indicates original feature dimension.
[…]
1 73.6 62.3 66.6 39.1
(g) Shared experts. Not sharing experts yields better performance.
[…]
(h) Number of experts. More experts improve performance.
[…]
4 74.0 64.9 66.6 38.1
[…]
(i) Batch size. Larger batchsize improves performance.
We investigate the impact of MoE design choices and hyper-parameters choices in indoor cross-domain 3D semantic segmentation. All experiments are conducted with ScanNet [21], Structured3D [88] and S3DIS [2] and evaluate without providing dataset label on seen dataset and also evaluate on unseen dataset: Matterport3D [12]. Results are presented in Tab. 3.
Auxiliary load-balancing loss. The auxiliary load-balancing loss is designed to encourage uniform expert usage and prevent expert collapse [54, 76, 43]. We ablate its effect by training models with varying strengths of this loss in Tab. 3. Interestingly, we find that removing the auxiliary loss entirely leads to consistently better performance across domains. In contrast, stronger auxiliary losses result in significant degradation. We hypothesize that this is due to the inherent imbalance in the distribution of samples across domains in 3D point cloud datasets—a phenomenon also observed by ChartMoE [82].
Effect of activated experts. Tab. 3 compares the performance of configurations with 1, 2 and 3 activated experts, finding diminishing return when increasing number of activated experts. We find that activating two experts yields the best performance across domains. This analysis suggests that merely increasing the number of activated experts does not guarantee improved performance.
MoE position. We ablate the placement of MoE layers to understand where expert specialization is most effective. Empirically, we observe greater expert diversity when MoE is applied in the decoder compared to the encoder. In Tab. 3, we compare two configurations: applying MoE layers only in the decoder, versus applying them in both the encoder and decoder. Results show that applying MoE to both stages consistently yields better performance, suggesting that expert specialization benefits from exposure to both encoder and decoder.
Effect of scaling model with MoE. Tab. 3 compares Point-MoE with 4 and 8 experts with activated experts fixed at 2. Increasing the number of experts from 4 to 8 leads to improved performance on seen domains, and yields a modest gain of 0.4% on the unseen domain. Tab. 3 ablates the expert width by varying the intermediate size of the MLP (e.g., 1H, 2H). The results show a mixed trend: wider experts improve performance on ScanNet and Matterport3D, but degrade results on Structured3D and S3DIS. Finally, we investigate the effect of adding shared experts [63, 72, 22], which are applied universally without gating. We find that disabling shared experts consistently leads to better accuracy and domain specialization across all benchmarks, as reported in Tab. 3.
Ablation of hyper-parameters choices. We provide justification for our selected hyperparameters. In Tab. 3, BatchNorm [34] consistently outperforms other normalization methods [4, 85] in the cross-domain setting. Tab. 3 shows that ReLU [1] slightly outperforms SiLU [27]. Additionally, Tab. 3 demonstrates that importance of batch size in MoE training. Specifically, the batch size must exceed the number of experts to avoid performance degradation. In general, larger batch sizes lead to improved performance across domains.
4.4 Analysis of MoE
Figure 2: Training Loss and Validation mIoU over time. Left: training loss curves for joint training across multiple domains. Right: validation mIoU on indoor datasets. Point-MoE-L achieves competitive performance to PPT, and both significantly outperform the baseline PTv3-L.
Training efficiency and validation mIoU. Fig. 2 shows the training loss and validation mIoU curves for four models: the baseline PTv3-L, its improved variant PTv3-Mix-LN (which incorporates mixed-domain batches and LayerNorm), PPT-L [80], and our proposed Point-MoE-L. All models are trained from scratch across multiple domains. Point-MoE-L converges faster and achieves strong validation mIoU without using explicit dataset labels, matching or exceeding the performance of PPT-L trained with ground-truth domain labels. While all models reach similar training loss, only Point-MoE-L and PPT-L generalize effectively—reinforcing that low training loss is not indicative of strong cross-domain performance. On ScanNet, Structured3D, Matterport3D, and nuScenes, Point-MoE-L shows consistent improvement and avoids early plateaus seen in PPT-L, especially on Structured3D, suggesting stronger long-term learning. PTv3-L fails to generalize and exhibits unstable validation curves. PTv3-Mix-LN improves stability over PTv3-L significantly, but still underperforms Point-MoE-L and PPT-L.
Figure 3: Expert Choices Visualization. Different color represents a different expert. (a) shows a specific expert focusing on edges across the scene, revealing spatially coherent routing. (b) and (c) highlight expert selections in indoor scenes, where semantically related regions (e.g., chair, desk) are consistently routed to the same expert. (d) displays an outdoor scene with sparse LiDAR points, yet the model still demonstrates meaningful routing despite limited geometry.
Expert choice visualization. Fig. 3 showcases expert assignments for a validation scene at selected layers. In (a), we observe that early encoder layers rely heavily on geometric cues for routing. For instance, green experts are consistently activated along object boundaries such as the edges of desks and chairs, while red experts dominate flat surfaces. In (b) and (c), the decoder layers exhibit more semantically meaningful expert selection—likely due to their proximity to the loss function—with distinct experts attending to objects like desks, chairs, floors, and walls. In (d), we examine an outdoor scene with sparse LiDAR data. Despite limited geometric structure, the model still organizes routing meaningfully: nearby points are routed to blue experts, while farther points activate red experts. We include more visualizations in the appendix for completeness. We also note occasional visual artifacts where isolated points are assigned different experts than their neighbors, which may be related to PTv3’s architectural choices such as point serialization or positional encoding.
Figure 4: Expert routing behavior across domains. The left panel shows the most frequently used expert paths through each encoder (E-) and decoder (D-) layer, with feature channel sizes marked in parentheses, illustrating the variation in expert selection for tokens from different datasets. The router tends to assign experts based on the input domain, particularly in the decoder MoE layers. To quantify this, we compute the Jensen-Shannon Divergence (JSD) between the expert selection distributions of different datasets at each MoE layer.
Token pathways. To understand how Point-MoE adapts to diverse domains, we analyze expert routing behavior at the token level. Specifically, we track the top-1 expert assignment for each token across all MoE layers and construct full routing trajectories from the final layer back to the first. We then identify the top 100 most frequent expert paths based on their occurrence across all tokens. As shown in Fig. 4, encoder expert paths are substantially less diverse than those in the decoder, indicating that encoder layers perform more domain-agnostic processing. Interestingly, we observe a sparse routing pattern in deeper encoder layers. This may be attributed not to feature reuse, but to the U-Net-style design where features are spatially deep but token sparsity increases, resulting in reduced variability in routing decisions.
When examining domain-level trends, we find that certain dataset pairs—such as SemanticKITTI and nuScenes or ScanNet and Structured3D—share similar expert pathways, suggesting that Point-MoE implicitly clusters domains with related geometric or semantic structures. To quantify these observations, we compute the Jensen-Shannon Divergence (JSD) [46] between expert selection distributions across datasets at each MoE layer. JSD [46] is an entropy-based measure of divergence between expert routing distributions across datasets, weighted by their token proportions; its formal definition is provided in the supplementary. As shown in Fig. 4 (right), decoder layers exhibit significantly higher JSD, indicating stronger domain-specific specialization. Several encoder layers also display nontrivial JSD, underscoring the benefit of placing MoE throughout the network.
5 Discussions
This work embraces the “bitter lesson” [66] in AI: scalable generalization emerges from flexible architectures trained on diverse data, rather than hand-engineered domain priors [68]. We introduce Point-MoE, a mixture-of-experts architecture for 3D understanding that dynamically routes tokens to specialized experts across both indoor and outdoor domains. Our experiments demonstrate that Point-MoE not only achieves competitive performance on seen datasets but, crucially, also generalizes effectively to novel domains without explicit domain labels or supervision. Through careful analysis of expert routing behavior, we show that the model self-organizes around both geometric and semantic structures, adapting its computation pathways to the nature of the input. These results highlight the potential of sparse, modular computation for robust and scalable 3D scene understanding. We provide more analysis of Point-MoE and discuss our limitations in the supplementary material.
6 Acknowledgement
The authors gratefully acknowledge Research Computing at the University of Virginia for providing the computational resources and technical support that made the results in this work possible. Zezhou Cheng acknowledges the Adobe Research Gift, as well as GPU support provided by the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program and the National Artificial Intelligence Research Resource (NAIRR) Pilot program.
Appendix
For a thorough understanding of our Point-MoE, we have compiled a detailed Appendix. The table of contents below offers a quick overview and will guide to specific sections of interest.
[…]
Appendix A Limitations, Licenses and Risks
A.1 Limitations
Our work is limited by computational resource constraints, which restrict the scale of both the Point-MoE architecture and the training datasets. While our results already demonstrate strong generalization across diverse 3D domains, larger models and a broader training corpus could further enhance performance. Additionally, our expert routing relies solely on the standard gating loss without auxiliary regularization for load balancing or diversity, which may lead to suboptimal expert specialization. Finally, the dynamic routing mechanism incurs nontrivial communication and synchronization overhead compared to dense backbones like Point Transformer V3, highlighting the need for more efficient expert parallelism or conditional computation in future work.
A.2 Artifacts and License
We report a list of licenses for all datasets and code used in our experiment in Tab. 4. We strictly follow all the datasets and code licenses and limit the scope of these datasets and code to academic research only.
Table 4: License information for the scientific artifacts used.
Data Sources URL License
ScanNet [21] Link ScanNet Terms of Use
Structured3D [88] Link Structured3D Terms of Use
S3DIS [2] Link CC BY-NC 4.0
Matterport3D [12] Link Matterport3D License
nuScenes [10] Link CC BY-NC-SA 4.0
SemanticKITTI [6] Link CC BY-NC-SA 3.0
Waymo Open Dataset [73] Link Waymo Research License
Software Code URL License
Pointcept [20] Link MIT License
A.3 Ethical concerns and Risks
This study does not involve human subjects or annotators. All experiments are performed on publicly released 3D benchmarks—ScanNet [21], Structured3D [88], S3DIS [2], Matterport3D [12], SemanticKITTI [6], nuScenes [10], and Waymo [65]—each distributed under their respective licences. Although these datasets are widely used, they may still reflect geographic, socioeconomic, or sensor–specific biases, which Point-MoE could inherit or amplify when deployed in real-world applications. We therefore encourage future work to develop debiasing techniques, to curate more balanced point-cloud corpora, and to incorporate fairness-oriented evaluation metrics so that cross-domain 3D models can serve all communities responsibly.
Appendix B Additional Experiments
B.1 Experiment Settings
Data. We compare a diverse set of 3D semantic segmentation datasets spanning both indoor and outdoor environments. ScanNet [21] is a large-scale indoor RGB-D dataset reconstructed from real-world scans of apartments and offices. S3DIS [2] contains real indoor scans of commercial buildings, providing richly annotated scenes with diverse architectural structures. Structured3D [88] is a synthetic dataset based on photo-realistic rendered scenes from CAD models, offering clean labels and dense annotations. Matterport3D [12] consists of real RGB-D scans captured in a wide range of residential environments, covering multiple floors and room types.
For outdoor settings, SemanticKITTI [6] provides point clouds captured by a 64-beam LiDAR mounted on a car, annotated for semantic understanding of dynamic street scenes. Waymo Open Dataset [73] extends this setup with more diverse driving environments and higher annotation quality, collected using a custom autonomous vehicle sensor suite. nuScenes [10] offers multimodal driving data, including LiDAR scans, captured in urban environments with complex traffic scenarios and longer temporal contexts. We include detailed dataset statistic in Tab. 5.
Furthermore, we provide detailed per-class comparisons on semantic classes for both indoor and outdoor datasets, as shown in Fig. 8 and Fig. 8.
Table 5: Overview of 3D Semantic Segmentation datasets differences. L, W, and H represent the length, width, and height. For all datasets, we use the raw point clouds to obtain the point ranges.
Dataset Sensor Point Range (m) Scene Type
ScanNet [21] Reconstructed L=[-2.13, 17.12], W=[-1.46, 18.19], H=[-0.38, 6.97] indoor
S3DIS [2] Reconstructed L=[-37.93, 29.93], W=[-26.08, 46.06], H=[-2.65, 6.58] indoor
Structured3D [88] Synthetic L=[-53.70, 52.05], W=[-82.07, 53.43], H=[-0.78, 15.49] indoor
Matterport3D [12] RGB-D Camera L=[-34.99, 72.17], W=[-57.85, 75.65], H=[-6.79, 10.21] indoor
SemanticKITTI [6] 64-beam LiDAR L=[-81.15, 80.00], W=[-80.00, 79.99], H=[-32.49, 3.51] outdoor
Waymo [73] 64-beam LiDAR L=[-74.98, 76.53], W=[-75.00, 75.00], H=[-18.36, 6.78] outdoor
nuScenes [10] 32-beam LiDAR L=[-105.17, 105.25], W=[-105.30, 105.17], H=[-53.42, 20.01] outdoor
Training. Our default joint training strategy and Point-MoE-L configs are summarized in Tab. 8. We experiment with two batching schemes: (1) single batch, where each iteration samples points exclusively from one dataset, and (2) mix batch, where samples are drawn from multiple datasets with equal probabilities. We find that larger batch sizes lead to more stable expert assignment and better convergence behavior in the Mixture-of-Experts architecture. All models are trained from scratch on the joint multi-domain datasets without any fine-tuning stage.
Table 6: Categorical settings of Indoor Datasets.
Dataset #C
wall
floor
cabinet
bed
chair
sofa
table
door
window
bookshelf
bookcase
picture
counter
desk
shelves
curtain
dresser
pillow
mirror
ceiling
refrigerator
television
shower curtain
nightstand
toilet
sink
lamp
bathtub
garbagebin
board
beam
column
clutter
otherstructure
otherfurniture
otherprop
other
Structured3D 25
ScanNet 20
S3DIS 13
Matterport3D 21
Table 7: Categorical settings of Outdoor Datasets.
Dataset #C
car
bicycle
motorcycle
truck
other-vehicle
person
bicyclist
motorcyclist
road
parking
sidewalk
other-ground
building
fence
vegetation
trunk
terrain
pole
traffic-sign
barrier
bus
construction_vehicle
pedestrian
traffic_cone
trailer
driveable_surface
other_flat
manmade
other vehicle
sign
traffic light
construction cone
tree trunk
curb
lane marker
other ground
walkable
SemanticKITTI 19
nuScenes 16
Waymo 22
Table 8: Training settings. H indicates feature dimension.
Joint-training Indoor Joint-training Indoor-Outdoor Point-MoE-L
Config Value Config Value Config Value
optimizer AdamW optimizer AdamW # experts 8
scheduler OneCycleLR scheduler OneCycleLR topk 2
learning rate 0.005 learning rate 0.005 shared expert 0
weight decay 0.05 weight decay 0.05 activation func. ReLU
batch size 6 batch size 16 auxiliary loss 0
datasets ScanNet (1), S3DIS (1), Struct.3D (1) datasets ScanNet (1), S3DIS (1), Struct.3D (1), nuScenes (1), SemanticKITTI (1) normalization BatchNorm
domain-aware False
domain rand. False
iters 120k iters 160k expert capacity 2H
(a) Detector trained on indoor domains
(b) Detector trained on indoor-outdoor domains
Figure 5: Domain Classification Confusion Matrices. (a) Confusion matrix for a domain detector trained only on indoor datasets: Structured3D [88], ScanNet [21], and S3DIS [2]. (b) Confusion matrix for a detector trained on both indoor and outdoor datasets: Structured3D [88], ScanNet [21], S3DIS [2], SemanticKITTI [6], and nuScenes [10]. In both cases, we evaluate generalization to unseen domains. Values are normalized per predicted domain (row-wise).
B.2 Additional Results and Analysis
Domain Detector. To enable domain-aware models at inference time—where ground-truth dataset labels are unavailable—we train a lightweight classifier to predict the domain of an input point cloud. This domain detector serves a purely functional role: it allows domain-conditioned variants of Point-MoE and other baselines to operate in realistic deployment scenarios where users may supply point clouds from unknown or mixed sources. During training, we observe that strong data augmentations (e.g., grid dropout or random cropping) reduce the classifier’s reliability by masking domain-specific cues. To preserve dataset-level characteristics critical for this task, we disable most augmentations and apply only minimal point cropping for efficiency. The domain detector converges quickly, with accuracy stabilizing after just a few epochs. As shown in Fig. 5, it achieves high classification accuracy—99.2% for indoor-only and 95.5% for indoor-outdoor settings—making it suitable for practical use.
The high accuracy of the domain detector suggests that 3D datasets exhibit strong dataset-specific patterns, even when covering similar semantic categories. While we do not perform a detailed analysis of the underlying factors, differences in data acquisition, such as synthetic rendering, RGB-D reconstruction, or LiDAR scanning, are likely to introduce distinct geometric and distributional biases. These biases are implicitly learnable and can be exploited by simple classifiers. For conventional models, such domain discrepancies often necessitate careful data curation or normalization strategies to improve generalization. However, our findings with Point-MoE offer a different perspective: by explicitly modeling distributional variation through expert specialization, the architecture may naturally accommodate such domain shifts without requiring aggressive data homogenization.
Does Point-MoE work on single dataset?
Table 9: Performance of Point-MoE vs. PTv3 on individual datasets. We evaluate both models independently on each dataset to examine whether Point-MoE offers benefits even without cross-domain training. Point-MoE performs comparably to PTv3.
Method ScanNet Structured3D Matterport3D nuScenes Waymo Average
PTv3 75.01 79.37 53.89 77.16 70.20 71.13
Point-MoE 75.63 79.29 53.47 78.24 69.50 71.23
We evaluate Point-MoE without cross-domain supervision by training and testing independently on three indoor datasets (ScanNet [21], Structured3D [88], Matterport3D [12]) and two outdoor datasets (nuScenes [10], Waymo [65]). On ScanNet and nuScenes, Point-MoE outperforms PTv3 by +0.62 and +1.08, respectively. On Structured3D (evaluated on its test split), performance is nearly identical (–0.08). On Matterport3D (test split) and Waymo (validation split), Point-MoE underperforms PTv3 by –0.42 and –0.70, respectively. On average, Point-MoE achieves 71.23 vs. 71.13 for PTv3 (+0.10). These results suggest that expert specialization can emerge in single-domain settings, but the overall improvement is small and not consistent across datasets. We consider the findings inconclusive.
We hypothesize that Point-MoE may require different configurations or training strategies when applied to a single dataset. In particular, the current gating and load balancing mechanisms are optimized for multi-domain settings, where diverse token distributions benefit from expert diversity. In contrast, single-domain training may benefit from fewer experts, a larger top-$k$, or stronger regularization to avoid expert under-utilization.
Table 10: Joint Training Results on Indoor 3D Semantic Segmentation without Precise Evaluator. Each model is trained jointly on three indoor datasets and evaluated under a domain-agnostic setting without relying on ground-truth domain labels during inference. We report performance when models are required to predict across all domains. To evaluate generalization with a single “generic“ label, results marked (G) use this label for both seen and unseen domains. Unmarked results rely on domain detector predictions..
Methods Params Activated Mix Batch Layer. ScanNet Structured3D S3DIS Matterport3D
Val Val Val Val
Single-dataset training
PTv3-ScanNet [78] 46M 46M 74.4 19.4 8.5 37.9
PTv3-Structured3D [78] 46M 46M 20.0 78.0 7.0 19.5
PTv3-S3DIS [78] 46M 46M 4.9 3.4 66.1 5.2
PTv3-Matterport3D [78] 46M 46M 52.7 23.0 9.5 53.6
Domain-agnostic joint training
[…]
Domain-aware joint training
[…]
Table 11: Cross-Domain 3D Semantic Segmentation on Indoor and Outdoor Scenes without Precise Evaluator. Models are trained jointly on both indoor and outdoor datasets under domain-agnostic supervision and domain-aware joint training settings. (G): Using the “generic“ label during evaluation (see definition above).
Methods Params Activated Mix Layer. ScanNet Structured3D S3DIS nuScenes SemKITTI Matterport3D Waymo
Val Val Val Val Val Val Val
Domain-agnostic joint training
[…]
Domain-aware joint training
[…]
Table 12: Joint Training Performance on Indoor 3D Semantic Segmentation with Ground-Truth Domain Labels. We additionally evaluate models trained on indoor datasets using domain-aware supervision, where explicit ground-truth domain labels are provided during training to guide the learning process.
Methods Params Activated Mix Batch ScanNet Structured3D S3DIS
Val Val Test Val
Domain-aware joint training with G.T. labels
PPT-S [80, 74] 47M 47M 74.7 60.9 65.5 71.6
PPT-L [80, 74] 98M 98M 74.6 60.9 62.6 69.2
Point-MoE-S 59M 52M 76.7 66.7 72.3 71.9
Point-MoE-L 100M 60M 75.6 69.7 71.3 69.6
Point-MoE-G-S 59M 52M 75.7 69.1 70.8 68.1
Point-MoE-G-L 100M 60M 76.7 68.1 69.6 67.1
Table 13: Joint Training Performance on Indoor-Outdoor 3D Semantic Segmentation with Ground-Truth Domain Labels. We evaluate models trained on indoor-outdoor datasets using domain-aware supervision with ground truth labels.
Methods Params Activated Mix ScanNet Structured3D S3DIS nuScenes SemKITTI
Val Val Test Val Val Val
Domain-aware joint training with G.T. labels
PPT-L [80, 74] 98M 98M 76.7 67.2 66.0 71.2 72.6 66.9
Point-MoE-L 100M 60M 76.8 69.6 70.1 70.6 72.4 64.5
Point-MoE-G-L 100M 60M 77.0 74.9 73.2 71.6 73.0 64.3
B.3 Additional Main Results
Results without precise evaluator. The results reported in Tab. 1 and Tab. 2 are obtained using the precise evaluator provided by Pointcept [20]. The precise evaluator first splits each scene into multiple overlapping chunks, then performs model inference on each chunk individually. After inference, it aggregates predictions across chunks through a voting mechanism to refine the final output. We typically observe a gain of 2–4% in accuracy on each dataset using this method. While effective, the precise evaluator introduces additional runtime overhead during evaluation. Moreover, we note that this form of ensembling may hurt overall performance when the model’s predictions are unreliable in early chunks, as incorrect outputs are also accumulated in the final vote. To support convenient and reproducible benchmarking for the community, we include results with the precise evaluator in the main text and provide additional baseline results without it elsewhere in the paper.
From both Tab. 1 and Tab. 2, we observe that Point-MoE outperforms the baselines in most domains in both the indoor-only and indoor-outdoor joint training settings. These results are consistent with those obtained using the precise evaluator reported in the main text. We do not observe any contradictory trends, as the precise evaluator typically improves all methods by a similar margin of 2–4%. As such, our analysis and conclusions align closely with those presented in the main text. How Strong is Domain-Aware Point-MoE? In the main text, we compare domain-agnostic and domain-aware settings and observe that domain-aware Point-MoE does not consistently outperform its domain-agnostic counterpart. Here we want to offer discussions and analysis to explain the reason. One contributing factor is the reliance on the domain predictor as we detailed in sec. B.2. Although the domain predictor achieves high classification accuracy, even minor errors can lead to noticeable performance degradation—either by misrouting inputs to suboptimal experts in Point-MoE or by selecting incorrect normalization statistics in PPT. To better isolate this effect, we include results in Tab. 12 and Tab. 13, where we evaluate domain-aware Point-MoE using ground-truth domain labels.
In general, Point-MoE continues to outperform standard baselines in both settings. Moreover, we find that domain-aware Point-MoE with access to ground-truth labels achieves better performance on seen domains, while domain-agnostic Point-MoE trained with domain-randomized data generalizes better to unseen domains. Interestingly, we also observe that Point-MoE trained with generic (i.e., domain-agnostic) supervision performs well even on seen domains when exposed to more diverse training inputs. This suggests that domain-randomization during training may benefit performance across both seen and unseen domains in challenging settings. Finally, we note that domain-aware Point-MoE can yield stronger results when accurate domain labels are available at inference time.
Appendix C Additional Visualization and Discussions on Point-MoE
C.1 t-SNE
Figure 6: t-SNE Visualization. The first three columns illustrate feature similarities and differences across PTv3, PPT, and Point-MoE for both encoder and decoder layers. The rightmost column highlights how Point-MoE generalizes to unseen domains, with Matterport3D features aligning closely with the most relevant seen domains—demonstrating robust cross-domain generalization.
Fig. 6 presents t-SNE visualizations [69] of encoder and decoder features from PTv3, PPT [80], and Point-MoE. PTv3 exhibits no clear separation in either encoder or decoder spaces, indicating its inability to structure domain-specific representations. In contrast, PPT shows strongly separated clusters for each domain at both encoder and decoder levels, suggesting it overfits to domain-specific patterns. While this benefits seen domains, it may hinder generalization to unseen settings.
Point-MoE, however, reveals a more interesting behavior. Its encoder features remain entangled across domains—indicating shared representation learning—while its decoder learns to disentangle domain-specific structures. This division of labor suggests that Point-MoE effectively performs domain inference within the model, without requiring explicit domain labels. Most notably, in the unseen Matterport3D domain, Point-MoE maps samples to nearby clusters of the most relevant seen domains (e.g., ScanNet), enabling strong generalization. This implicit alignment underscores the model’s ability to organize and interpret new environments by leveraging expertise learned from diverse sources.
C.2 Semantic Concept Distribution over Experts
Expert Specialization Word Cloud. To visualize expert specialization, we collect the semantic predictions from each expert in the decoder layers of Point-MoE. During inference, we log which expert is selected for each point and record the corresponding semantic prediction. For each expert, we compute the frequency distribution of predicted classes across the entire validation set and display the most frequent class as a word cloud. Colors indicate the originating dataset of the prediction (e.g., ScanNet, S3DIS, etc.), providing insight into cross-domain consistency.
We observe several interesting patterns in the expert specialization behavior. (1) In each decoder layer, there is typically at least one expert with a strong preference for outdoor data. For example, Decoder 0_0 Expert 7, Decoder 1_1 Expert 1, and Decoder 2_1 Expert 2 predominantly handle outdoor scenes. (2) Generic or ambiguous semantic classes such as otherfurniture or other-ground never dominate the assignment of any expert, suggesting that these classes are more diffusely distributed and do not concentrate through the routing mechanism.
From the visualization in Fig. 7, we observe several notable trends. Specialization emerges clearly: within each decoder block, experts tend to focus on semantically coherent regions such as bed, bathtub, window, or trailer, suggesting that experts develop functional roles under shared supervision. In many blocks, multiple experts cover similar classes, yet differ in dataset association, indicating potential cross-domain generalization. For instance, certain experts across blocks frequently predict bed or toilet across multiple indoor datasets, hinting at robust feature preferences. We also observe signs of dataset sensitivity. Classes like trailer, parking, and terrain appear predominantly in SemanticKITTI and nuScenes, while sink, shower curtain, and dresser are more common in Structured3D. Comparing across decoder layers, earlier blocks display a broader mix of classes per expert, while deeper blocks exhibit more focused and object-specific specialization, reflecting an increasing semantic refinement through the decoding process.
These findings suggest that Point-MoE encourages structured representation learning, where each expert in a given layer implicitly captures recurring semantic or geometric patterns, even without explicit specialization objectives or layer-to-layer consistency.
Figure 7: Expert Specialization Word Cloud. All classes’ frequencies of selecting their top-1 expert are measured on the validation set. Colors indicate the originating dataset of the class (e.g., ScanNet, S3DIS, etc.), providing insight into cross-domain consistency.
Expert Probability Stack Visualization. To complement the word cloud analysis, we provide fine-grained views of expert behavior in Fig. 8, Fig. 9, Fig. 10 and Fig. 11. For each semantic class, we visualize the soft routing probabilities across all experts within a specific decoder block. Each horizontal bar corresponds to a class, segmented proportionally by the probability of top-1 expert assignment to each expert. This representation allows us to examine how responsibilities are distributed across experts for different categories.
We observe that certain semantic classes are strongly associated with a small number of experts. For example, bed, bathtub, and toilet are consistently routed to a narrow subset of experts, reflecting high specialization. In contrast, classes such as wall and vegetation are distributed across multiple experts, likely due to their spatial extent and ubiquity in scenes. Some experts show broad activation across diverse classes, while others are more selective, indicating a balance between general-purpose and task-specific routing. Several consistent patterns emerge in the decoder. In Decoder Layer 3 Block 0, both truck and trailer are primarily assigned to Expert 3, suggesting structural similarity. In Decoder Layer 3 Block 1, dresser, nightstand, and counter are all routed to Expert 2, reflecting shared visual or functional features. In Decoder Layer 0 Block 1, traffic cone and barrier exhibit nearly identical routing distributions, likely due to their geometric similarity and outdoor context. These observations confirm that the model learns structured yet flexible routing strategies, shaped by both semantic identity and spatial context.
Figure 8: Expert Choices Visualization for Decoder Layer 0.
Figure 9: Expert Choices Visualization for Decoder Layer 1.
Figure 10: Expert Choices Visualization for Decoder Layer 2.
Figure 11: Expert Choices Visualization for Decoder Layer 3.
C.3 Expert Evolution.
Figure 12: JSD Dynamics During Training. We visualize the Jensen-Shannon Divergence (JSD) between expert assignment distributions across datasets to study how expert specialization emerges over time. The plots illustrate JSD dynamics across training iterations for both encoder and decoder layers.
In Fig. 12, we track the Jensen-Shannon Divergence (JSD) between expert assignment distributions across different datasets every 10k training iterations to analyze how expert specialization evolves over time. Let $P_{1},P_{2},\dots,P_{n}$ denote the expert routing distributions for each of the $n$ datasets, where each $P_{i}$ is a categorical distribution over experts. Specifically, let $P_{i}(e)$ denote the fraction of tokens from dataset $i$ that are routed to expert $e$. Let $\pi_{1},\pi_{2},\dots,\pi_{n}$ be the corresponding weights assigned to each dataset, satisfying $\sum_{i=1}^{n}\pi_{i}=1$. These weights reflect the proportion of tokens contributed by each dataset. The Jensen-Shannon Divergence is defined as:
$\text{JSD}_{\pi}(P_{1},P_{2},\dots,P_{n})=H\mathopen{}\mathclose{{}\left(\sum_{i=1}^{n}\pi_{i}P_{i}}\right)-\sum_{i=1}^{n}\pi_{i}H(P_{i}),$
where $H(P)=-\sum_{x}P(x)\log P(x)$ denotes the Shannon entropy. This formulation captures the degree of divergence between datasets in terms of how they utilize the experts, providing a measure of specialization and separation over time.
In the early stages (10k–30k), JSD values remain low across most experts, indicating that the routing behavior is largely uniform across datasets and experts have not yet developed distinct roles. Between 40k and 60k iterations, we begin to observe divergence among certain experts, marking the onset of specialization. This trend continues and becomes more pronounced by 80k iterations, where several experts display increasingly distinct assignment distributions. By 100k iterations, the JSD curves stabilize, suggesting that expert roles have converged. Some experts diverge earlier and more sharply, indicating faster specialization, while others evolve more gradually. Overall, this analysis confirms that expert diversity in Point-MoE is not predefined but emerges progressively through training.
In the encoder, we observe that some experts maintain JSD values close to zero throughout training, suggesting that their routing behavior remains largely random and undifferentiated. This observation aligns with Fig. 4, where certain encoder experts do not exhibit clear dataset- or task-specific specialization. This raises the possibility that not all encoder layers require MoE modules, or that additional regularization or training signals may be needed to encourage expert differentiation in the encoder. Nonetheless, we do observe that some encoder experts do specialize over time, indicating that meaningful expert roles can still emerge under the right conditions.
In the decoder, we observe that all experts exhibit high JSD values, indicating that their routing behavior becomes highly specialized and dataset-dependent. Additional visualizations of expert assignments can be found in Sec. C.4, further illustrating this specialization.
C.4 More Expert Choice Visualization
Fig. 13, Fig. 14, Fig. 15, Fig. 16, Fig. 17, Fig. 18, and Fig. 19 showcase expert assignments for a validation scene across all layers. Due to the hierarchical structure of the PTv3 backbone, we observe denser point cloud representations at the input and output stages, with sparser intermediate features. This design leads to visibly varying point densities and detail levels across the network. We also observe block-like artifacts in some visualizations, which are likely attributed to PTv3’s point serialization strategy. In the encoder, certain experts occasionally exhibit collapsed or spatially clustered behavior—a phenomenon we refer to as “expert clasping.” A deeper investigation into this behavior is left for future work. In the decoder layers, expert assignments closely resemble semantic segmentation outputs, even though the visualizations reflect only expert routing decisions rather than predicted labels. This suggests that experts naturally specialize in semantically coherent regions, despite the absence of explicit supervision guiding their roles.
In Fig. 13, we observe that early encoder layers primarily rely on geometric cues for expert routing. For example, certain experts—such as the green one—are consistently activated along object boundaries, including the edges of desks and chairs, while others tend to dominate on flat or horizontal surfaces. In the decoder layers, expert assignments become more semantically meaningful, with different experts focusing on specific object categories such as desks, chairs, walls, and floors.
In Fig. 14, we observe a similar overall pattern to ScanNet [21]. In the encoder, expert assignments are primarily driven by local geometric structure, with clear attention to furniture boundaries and wall intersections. Notably, edge-aware routing persists deeper into the encoder—for example, in encoder layer 2 block 0, certain experts remain focused on object contours. In the decoder, expert activation becomes increasingly aligned with semantic regions, such as tables, chairs, and structural elements. Despite the architectural and layout differences from ScanNet, the expert routing in S3DIS remains semantically consistent across the scene.
Fig. 15 illustrates expert routing on Structured3D [88], a synthetic dataset. Compared to real-world datasets, the expert assignment patterns appear sharper and more consistent, likely due to the absence of sensor noise and reconstruction artifacts. Some experts are clearly focused on flat regions, such as beds or floor. In the decoder layers, expert specialization aligns closely with object instances, and the separation across semantic regions is more distinct and well-defined, highlighting the benefits of clean, synthetic geometry for expert differentiation.
Fig. 16 presents expert routing on Matterport3D [12], which contains complex residential environments. In the early encoder layers, we observe clear expert usage patterns—for instance, a distinct separation between vertical and horizontal surfaces. In the decoder layers, expert assignments become more semantically meaningful, with experts specializing in large structural elements such as floors and ceilings, as well as common household objects like counters, sofas, and chairs. Notably, Matterport3D is not included in the training set, so this behavior reflects Point-MoE’s ability to generalize expert specialization to unseen domains.
In Fig. 17, we examine an outdoor scene from SemanticKITTI [6], characterized by sparse LiDAR data. Despite the limited geometric density, the model still exhibits meaningful expert routing: nearby points tend to be assigned to certain experts, while distant points activate others. Although the transition from geometry-driven to semantically informed expert assignment remains observable, it appears noisier due to the inherent sparsity and irregularity of the input point cloud.
Fig. 18 shows expert routing on nuScenes [10], an outdoor dataset with lower-resolution LiDAR data. In the encoder layers, expert assignments appear more fragmented, likely due to the reduced point density. Nonetheless, we observe that closer points are often routed to specific experts, while distant points are handled by others, indicating a spatially aware routing pattern. In the decoder layers, expert assignments become more semantically coherent, with specialization emerging for vehicle bodies, building facades, and road surfaces based on point position and distance.
Fig. 19 presents expert routing on Waymo [65], an unseen dataset characterized by 64-beam LiDAR scans. In the encoder, experts exhibit structured activation patterns even in early layers, with clear delineation between flat surfaces and protruding objects. In the decoder layers, expert routing becomes more semantically aligned, with experts consistently activated over road surfaces, vehicle bodies, and trees. The assigned regions remain compact and well-separated.
Figure 13: Point-MoE’s Expert Choice Visualization on ScanNet. Each point is colored by its assigned expert across encoder and decoder blocks. Early layers show structured expert usage, while deeper layers exhibit more mixed routing, reflecting the evolving representation dynamics throughout the network.
Figure 14: Point-MoE’s Expert Choice Visualization on S3DIS. Each point is color-coded based on the expert it was routed to across both encoder and decoder blocks. The earlier layers demonstrate clear structure in expert assignments, whereas the later layers reveal more diverse routing patterns, indicating evolving feature representations.
Figure 15: Point-MoE’s Expert Choice Visualization on Structured3D. Points are colored by their designated experts in the encoder and decoder. While the initial layers display consistent expert allocation, the deeper layers show a blend of routing decisions, highlighting the network’s shifting representational strategy.
Figure 16: Point-MoE’s Expert Choice Visualization on Matterport3D. Color represents the expert assigned to each point throughout the encoder and decoder. Structured expert usage is visible in early stages, transitioning to more varied routing in deeper layers, signaling changes in how features are processed.
Figure 17: Point-MoE’s Expert Choice Visualization on SemanticKITTI. The expert responsible for each point, shown through color, spans the encoder and decoder blocks. Initially, expert usage is highly organized, but becomes more dispersed in later layers, capturing the transformation in representation over depth.
Figure 18: Point-MoE’s Expert Choice Visualization on NuScene. Each point is colored according to its corresponding expert across the network’s layers. Early encoder and decoder blocks exhibit distinct expert patterns, while deeper layers demonstrate a more distributed and dynamic routing behavior.
Figure 19: Point-MoE’s Expert Choice Visualization on Waymo. Color indicates the expert each point is assigned to across encoder and decoder stages. Early network layers maintain structured expert routes, but this structure relaxes in deeper layers, illustrating evolving internal representations.

### v2

@@
-Point-MoE: Towards Cross-Domain Generalization in 3D Semantic Segmentation via Mixture-of-Experts
+Point-MoE: Large-Scale Multi-Dataset Training with Mixture-of-Experts for 3D Semantic Segmentation
 Abstract
-While scaling laws have transformed natural language processing and computer vision, 3D point cloud understanding has yet to reach that stage.
-This can be attributed to both the comparatively smaller scale of 3D datasets, as well as the disparate sources of the data itself.
-Point clouds are captured by diverse sensors (e.g., depth cameras, LiDAR) across varied domains (e.g., indoor, outdoor), each introducing unique scanning patterns, sampling densities, and semantic biases.
-Such domain heterogeneity poses a major barrier towards training unified models at scale, especially under the realistic constraint that domain labels are typically inaccessible at inference time.
-In this work, we propose Point-MoE, a Mixture-of-Experts architecture designed to enable large-scale, cross-domain generalization in 3D perception.
-We show that standard point cloud backbones degrade significantly in performance when trained on mixed-domain data, whereas Point-MoE with a simple top-$k$ routing strategy can automatically specialize experts, even without access to domain labels.
-Our experiments demonstrate that Point-MoE not only outperforms strong multi-domain baselines but also generalizes better to unseen domains.
-This work highlights a scalable path forward for 3D understanding: letting the model discover structure in diverse 3D data, rather than imposing it via manual curation or domain supervision.
+While massively scaling both data and models have become central in NLP and 2D vision, their benefits for 3D point cloud understanding remain limited.
+We study the initial step of scaling 3D point cloud understanding under a realistic regime: large-scale multi-dataset joint training for 3D semantic segmentation, with no dataset labels available at training or inference time.
+Point clouds arise from a wide range of sensors (e.g., depth cameras, LiDAR) and scenes (e.g., indoor, outdoor), yielding heterogeneous scanning patterns, sampling densities, and semantic biases; naively mixing such datasets degrades standard models.
+Therefore, we introduce Point-MoE, a Mixture-of-Experts design that expands model capacity through sparsely activated expert MLPs and a lightweight top-$k$ router, allowing tokens to select specialized experts without requiring dataset supervision.
+Trained jointly on a diverse mix of indoor and outdoor datasets, and evaluated on seen datasets as well as in zero-shot settings, Point-MoE outperforms prior methods without using dataset labels for either training or inference.
+This outlines a scalable path for 3D perception: letting the model discover structure in heterogeneous 3D data rather than imposing it via manual curation or dataset-specific heuristics.
 1 Introduction
-In both language and vision, we have witnessed remarkable advances driven by two trends: the aggregation of massive heterogeneous datasets [59, 64, 5, 57, 9, 89], and the deployment of models of ever-increasing capacity trained on them in a unified manner [8, 18, 55, 39, 25, 49].
-These models succeed not because they were crafted for a specific domain, but because they were trained to learn underlying patterns across a wide variety of datasets.
-However, 3D point cloud understanding has yet to follow the same trajectory of progress, despite its wide applications in robotics[13, 50, 84], autonomous systems[15, 14, 44], augmented reality[90, 75], and embodied intelligence[77, 71].
-Taking the 3D semantic segmentation task as an example, although diverse 3D datasets exist—such as ScanNet [21], SemanticKITTI [6], nuScenes [11], and ASE [3]—each covers only a narrow subset of the real-world diversity in 3D assets.
-Point clouds are generated from a variety of pipelines: RGB-D sensors [12], terrestrial and airborne LiDAR [33, 51], and multi-view stereo [83].
-Each source brings its own characteristics—differing point densities, coverage patterns, reconstruction artifacts, and semantic biases (Fig.
-1 left).
-The result is a 3D ecosystem of rich, yet siloed data sources, where models trained in one silo often fail to generalize across others.
-A natural response to address this challenge is to train across such silos.
-For example, one might naively train the state-of-the-art Point Transformer V3 (PTv3) [78] (Fig.
-1a) on mixed-domain 3D datasets.
-However, such models struggle to reconcile heterogeneous characteristics of the data, even with increased network capacity (Tab.
+In both language and vision, we have witnessed remarkable advances driven by two trends: the aggregation of massive, heterogeneous datasets [61, 66, 4, 58, 8, 97], and the deployment of ever-larger models trained on them in a unified manner [7, 16, 57, 36, 21, 51].
+These models succeed not because they are crafted for a specific dataset, but because they learn underlying regularities across multiple datasets.
+However, 3D point cloud understanding has not yet followed the same trajectory, despite its wide applications in robotics [11, 53, 90], autonomous systems [13, 12, 43], augmented reality [6, 77], and embodied intelligence [79, 72].
+Taking 3D semantic segmentation as an example, although diverse 3D datasets exist—such as ScanNet [19], SemanticKITTI [5], nuScenes [9], and ASE [2]—each covers only a narrow slice of real-world variation.
+Point clouds arise from diverse pipelines (RGB-D, LiDAR, multi-view stereo), each with their own distinct densities, artifacts, and semantic biases (Fig.
+1, left).
+The result is a rich yet siloed ecosystem of 3D models, where models trained on one domain often fail to generalize to others.
+A natural response is to train across silos, presumably resulting in a single set of model parameters that is effective across the spectrum of datasets.
+One might, for example, naively train the state-of-the-art Point Transformer V3 (PTv3) [81] (Fig.
+1a) on a mixture of datasets from multiple domains, e.g.
+indoor and outdoor.
+In practice, however, such models struggle to reconcile the heterogeneity of the data distribution even as capacity scales (Tab.
 1).
-To mitigate this, recent methods introduce domain-aware components.
+To mitigate this, recent methods have introduced dataset-aware components.
 Point Prompt Training (PPT) [82] (Fig.
-1b) modifies PTv3 by assigning domain-specific normalization layers.
-Similarly, One-for-All [74] achieves improved 3D detection performance by employing a lightweight classifier to infer dataset labels, then applying domain-specific adapters.
-While these methods show notable gains, they rely on access to domain labels during training and exhibit performance degradation on unseen domains.
-A likely cause is that adapting only the normalization parameters may be insufficient to fully capture the complex variations required for scalable, cross-domain generalization.
-In this work, we hypothesize that Mixture-of-Experts (MoE) architectures [35, 38] are well-suited for cross-domain joint training.
+1b) assigns dataset-specific normalization layers, while One-for-All [75] uses a lightweight dataset classifier with dataset-specific adapters.
+While effective, these methods rely on dataset labels, identifying which specific dataset an input sample originated from, during training and inference.
+Moreover, adapting only the normalization parameters may be insufficient for large-scale multi-dataset training.
+At deployment, point clouds arrive from mixed, unlabeled sources with unknown provenance, i.e.
+there is no oracle “dataset ID” available during inference.
+Accordingly, we study a realistic regime: large-scale multi-dataset joint training for 3D semantic segmentation with no dataset labels at inference.
+We hypothesize that Mixture-of-Experts (MoE) architectures [30, 34] are well-suited to this setting.
 We introduce Point-MoE, a sparse MoE model built upon PTv3 [81].
 As illustrated in Fig.
-1c, Point-MoE replaces each feed-forward layer in PTv3 with a MoE layer comprising of multiple expert networks (implemented as feed-forward MLPs) and a router that dynamically selects a sparse subset of experts for each input token.
-Our design is motivated by three key insights.
-First, MoE provides a natural inductive bias for handling diverse inputs by encouraging expert specialization and learning token-to-expert routing [17, 82].
-This allows a single model to be trained on mixed-domain data without requiring access to domain labels.
-Second, compared to normalization-based domain adaptation methods like PPT [80] and One-For-All [74], MoE offers a significantly larger and more expressive representation space, enabling better modeling of domain-specific variations.
-Third, sparse MoE architectures improve computational efficiency during both training and inference, and have emerged as a standard design for scaling large models [41, 28, 26, 58, 22].
-We extend these empirical benefits of MoE, previously established in language and vision, to the domain of 3D point cloud understanding, with a particular focus on 3D semantic segmentation.
-Figure 1: Overview of cross-domain training architectures.
-Point clouds exhibit diverse characteristics across domains.
-(a) Naively training Point Transformer V3 (PTv3) [78] on multi-domain data leads to degraded performance within each domain.
-(b) Point Prompt Training (PPT) [80] addresses this by adding domain-aware normalization parameters.
-(c) Our proposed Point-MoE tackles this challenge with Mixture-of-Experts (MoE), enabling dynamic expert specialization across domains.
-Domain labels can optionally be provided to the router to guide the expert selection.
-We conduct extensive experiments across a diverse mixture of 3D datasets, spanning indoor and outdoor scenes, synthetic and real data, and multiple sensor types.
-To comprehensively assess the effectiveness of Point-MoE, we evaluate it under both domain-agnostic (i.e., without access to domain labels) and domain-aware training regimes.
-Our results show that Point-MoE consistently adapts well to both settings, achieving strong generalization to unseen domains while outperforming competitive baselines, with improved training (Fig.
-2) and inference efficiency (Tab.
-1 and 2).
-We also conduct a thorough ablation study analyzing key design choices, including the number of experts, sparsity level, and the placement of MoE layers (Tab.
-3).
-Remarkably, even without domain supervision, Point-MoE learns to route input data meaningfully and to specialize experts organically—emerging from the data distribution itself (Fig.
-3 and 4).
-To the best of our knowledge, this work presents the first systematic exploration of the Mixture-of-Experts architecture for 3D point cloud understanding in the context of cross-domain generalization.
-These findings point toward a scalable path forward for 3D perception: instead of designing separate models for each dataset or domain, we should aim to develop unified systems that can adapt across the full spectrum of 3D data sources.
-In this regard, Point-MoE emerges as a promising and flexible framework for bridging domain gaps in large-scale 3D semantic segmentation.
+1c, Point-MoE replaces each attention projection layer in each PTv3 block with a MoE module composed of expert MLPs and a lightweight router that selects a sparse top-$k$ subset of experts per token.
+Our design is motivated by three insights.
+First, MoE encourages expert specialization and token-to-expert routing [15, 87], enabling a single model to learn from mixed datasets without access to dataset labels at training or test time.
+Second, compared to normalization-based adaptations such as PPT [82] and One-for-All [75], MoE offers far more expressive capacity to model dataset-specific variations.
+Third, sparse MoE architectures maintain computational efficiency during training and inference, and are a standard mechanism for scaling large models in NLP and 2D vision [40, 24, 22, 59, 20].
+We extend these benefits, well-established in language and 2D vision, to 3D point cloud understanding, focusing on 3D semantic segmentation.
+Figure 1: Overview of multi-datasets training architectures.
+Point clouds exhibit diverse characteristics across datasets.
+(a) Naively training Point Transformer V3 (PTv3) [81] on multi-datasets data leads to degraded performance within each domain.
+(b) Point Prompt Training (PPT) [82] addresses this by adding dataset-aware normalization parameters.
+(c) Our proposed Point-MoE tackles this challenge with Mixture-of-Experts (MoE), enabling dynamic expert specialization across datasets.
+We conduct extensive experiments across a diverse collection of 3D datasets spanning indoor and outdoor scenes, synthetic and real datasets, and multiple sensor modalities.
+We evaluate on both in-domain and zero-shot datasets, all without dataset labels.
+To comprehensively assess Point-MoE, we evaluate both in-domain (seen) datasets and out-of-domain (zero-shot) datasets, all without access to dataset labels.
+Our goal is to demonstrate that Point-MoE not only achieves strong performance on seen datasets but also generalizes robustly to unseen distributions, without relying on dataset-specific augmentation or supervision.
+Across both regimes, Point-MoE consistently outperforms competitive multi-dataset baselines.
+We further ablate the number of experts, sparsity level, and MoE placement (Tab.
+3), and find that Point-MoE learns coherent routing patterns and exhibits organic expert specialization driven by the underlying data distribution (Fig.
+234).
+To the best of our knowledge, this work provides the first systematic study of MoE for 3D point cloud understanding under the large-scale multi-dataset training regime.
+Our contributions are: (1) we focus on multi-dataset 3D semantic segmentation without domain labels, using a single model, and put forth Mixture-of-Experts as a natural solution to this task; (2) we carefully explore the MoE design space and provide extensive ablations that highlight the trade-offs and inform effective configurations; (3) we demonstrate that Point-MoE achieves state-of-the-art performance across seven datasets under this protocol while maintaining inference efficiency; and (4) we analyze token-level routing trajectories and expert usage, revealing encoder–decoder specialization patterns consistent with 3D geometry.
+Together, these findings point toward a scalable path for 3D perception: rather than building separate models for each dataset or domain, a single unified system can adapt across the spectrum of 3D data sources, leveraging scaling laws for data and compute [68].
 2 Related Work
@@
 Among them, Point Transformer V3 (PTv3) [81] represents the current state of the art, introducing a novel space-filling curve based serialization scheme, that transforms the unstructured 3D points into a 1D sequence, enabling efficient transformer-based modeling.
-A recent series of works have explored scaling such point cloud models to the multi-dataset setting using domain-aware training strategies.
-Point Prompt Training (PPT)  [80] demonstrates that using domain-specific normalization layers during pretraining improves 3D segmentation performance across multiple datasets.
-In 3D detection, One-for-All [74] adopts a sparse convolutional network with domain-specific partitions and normalization layers to boost multi-domain performance.
-Rather than relying on such domain-specific architectural branches [80, 74], we propose a unified sparse expert framework that learns to specialize dynamically across various domains, without requiring explicit dataset labels during training or inference.
+Recent works [32, 96, 39, 82, 83] scale point cloud understanding models to multi-dataset settings and show that large-scale point cloud data can improve performance on specific datasets: Sonata [83] explores multi-dataset self-supervised learning, but does not train across indoor and outdoor datasets.
+On the other hand, Point Prompt Training (PPT)  [82] demonstrates that using dataset-specific normalization layers during pretraining improves 3D semantic segmentation performance across multiple datasets along with dataset-specific fine-tuning.
+In 3D detection, One-for-All [75] adopts a sparse convolutional network with dataset-specific partitions and normalization layers to boost multi-datasets performance, and UniDet3D [37] achieves state-of-the-art performance with multi-dataset joint training by merging label spaces across datasets into unified taxonomy and building upon vanilla transformer [71] architecture.
+Rather than relying on dataset-specific preprocess or architectural branches [82, 75], we propose a unified sparse expert framework that learns to specialize dynamically across various datasets, without requiring explicit dataset labels during training or inference.
 Mixture-of-Experts.
-Mixture-of-Experts (MoE) [35, 38] aims to increase model capacity without a proportionate increase in compute costs by levereaging conditional computation – large portions of the model are only activated conditioned on certain input samples [60].
-In NLP, MoE has enabled scaling up models from hundreds of millions to hundreds of billions (and beyond) parameters: GShard scales to 600 B parameters [41], Switch Transformer to 1.6 T parameters [28], ST-MoE to 260 B parameters [91], and GLaM to 1.2 T parameters [26], all at roughly the same training cost as a dense model one-tenth the size.
-V-MoE [58] applies this recipe to vision, demonstrating competitive accuracy at 15 B parameters with batch-priority routing.
-More recently, Mixtral 8x7B [37] and DeepSeek models [22, 23, 24] all employ Mixture-of-Experts for large language model pretraining.
-Apart from being a powerful tool for scaling up model capacity, MoE is well-suited to training on data drawn from heterogeneous domains, because the experts naturally tend to self-organize on a per-domain basis.
-Early work shows gains in multi-source sentiment and POS tagging [30], domain-adaptive language modeling [32], automatic chart understanding [82], and multilingual translation [17].
-Mod-Squad [16] introduces a modular mixture-of-experts framework for multi-task learning, optimizing task-expert matching to balance cooperation and specialization across tasks.
-DAMEX [36] applies this to visual domains, while MoE-LLaVA [45], DeepSeek-VL [47], and DeepSeek-VL 2 [81] report similar benefits in vision-language models.
-To our knowledge, ours is the first attempt at extending MoE to address multi-dataset training and cross-domain generalization for 3D point clouds.
-3 Point-MoE for Cross-Domain Point Cloud Understanding
+Mixture-of-Experts (MoE) [30, 34] increases model capacity without proportional compute by activating only a subset of experts per input [62].
+This conditional computation has powered trillion-scale NLP models [40, 24, 98, 22, 33, 20, 46, 47] and demonstrated similar efficiency in vision [59].
+Beyond scaling, MoE is naturally suited for heterogeneous or multi-datasets data, as experts tend to self-organize by dataset or task.
+Prior work has shown improvements in multi-source sentiment analysis and POS tagging [26], dataset-adaptive language modeling [28], automatic chart understanding [87], and multilingual translation [15].
+Mod-Squad [14] introduced modular expert routing for multi-task learning, balancing cooperation and specialization.
+Similar trends are observed in vision and vision-language applications: DAMEX [31], MoE-LLaVA [44] and DeepSeek-VL [48, 85].
+To our knowledge, this is the first work to extend MoE to multi-dataset training in 3D point clouds.
+MoE3D [76] recently introduces a MoE formulation for 3D reconstruction.
+Domain generalization.
+Classical domain generalization (DG) trains on one or more labeled source domains and targets zero-shot performance on a specified target domain, typically in sensor- and semantic-ontology-homogeneous outdoor LiDAR settings [35, 60, 92, 84, 88, 86].
+We instead study large-scale joint training on heterogeneous datasets with naturally misaligned label spaces, evaluating a single model on both seen and zero-shot datasets.
+Unlike DG, our goal is a single unified model for heterogeneous multi-dataset learning rather than transfer to a fixed target domain, although DG methods remain complementary and could be layered onto Point-MoE as a natural extension.
+3 Point-MoE for Multi-Dataset Point Cloud Training
+3.1 Large-Scale Multi-Dataset Joint Training
 Task definition.
-Given a 3D scene represented as a point cloud $\mathcal{P}=\{p_{i}\}_{i=1}^{n}$, where each point $p_{i}\in\mathbb{R}^{3}$ denotes coordinates, semantic segmentation aims to predict a single class label $\hat{y}_{i}\in\mathcal{C}$ to each point $p_{i}$ from a fixed label set $\mathcal{C}=\{c_{1},c_{2},\ldots,c_{m}\}$.
-Multi-domain data are utilized during training.
-Let the domain containing $N$ point clouds be $\mathcal{D}=\mathopen{}\mathclose{{}\left\{\mathopen{}\mathclose{{}\left(p_{i},y_{i}}\right)_{i=1}^{n}}\right\}^{N}$.
-Then the complete training set across $d$ domains can be represented as: $\mathopen{}\mathclose{{}\left\{\mathcal{D}_{k}}\right\}_{k=1}^{d}$.
-The goal of multi-domain joint training is to learn a unified model over all domains that minimizes prediction error across $\mathopen{}\mathclose{{}\left\{\mathcal{D}_{k}}\right\}$, while also generalizing effectively to domains unseen during training.
-We adopt a minimal Mixture-of-Experts design that closely follows the standard architecture prevalent in the recent NLP literature.
-This simplicity allows us to leverage existing scalable Transformer-based MoE implementations with minimal modification.
-For the base model, we use Point Transformer V3 (PTv3) [78], a state-of-the-art architecture for point cloud understanding.
-Below we introduce the architecture of Point-MoE in detail.
-Mixture-of-Experts layer.
+Given a 3D scene represented as a point cloud $\mathcal{P}=\{p_{i}\}_{i=1}^{M}$, where each point $p_{i}\in\mathbb{R}^{3}$ denotes XYZ coordinates, semantic segmentation aims to predict a single class label $\hat{y}_{i}\in\mathcal{C}$ for each point $p_{i}$ from a fixed label set $\mathcal{C}=\{c_{1},c_{2},\ldots,c_{m}\}$.
+Predicted labels for the entire point cloud are compactly denoted as $\mathbf{\hat{y}}$.
+Similarly, we denote the corresponding ground-truth labels by $\mathbf{y}=[y_{1},y_{2},\ldots,y_{M}],y_{i}\in\mathcal{C},$ where $y_{i}$ is the true semantic label of point $p_{i}$.
+Let a dataset containing multiple point clouds be denoted as $\mathcal{D}=\mathopen{}\mathclose{{\left\{(\mathcal{P}_{t},\mathbf{y}_{t})}}\right\}_{t=1}^{|\mathcal{D}|}$.
+The complete training set across multiple datasets is then $\mathbb{D}=\mathopen{}\mathclose{{\left\{\mathcal{D}_{j}}}\right\}_{j=1}^{|\mathbb{D}|}$.
+The goal of our multi-dataset joint training is to learn a unified model (i.e.
+a single set of model parameters, without resorting to per-dataset fine-tuning) over all datasets that minimizes prediction error across $\mathbb{D}$, while also generalizing effectively to zero-shot datasets drawn from distributions unseen at training time.
+[…]
+We can formally contrast this to alternate multi-dataset objective formulations such as [82, 75], that learn an explicit set of dataset-specific parameters $\omega_{j}$, along with $\theta$: $\min_{\theta,\{\omega_{j}\}}\;\mathbb{E}_{\mathcal{D}_{j}\sim\mathbb{D}}\;\mathbb{E}_{(\mathcal{P},\,\mathbf{y})\sim\mathcal{D}_{j}}\mathopen{}\mathclose{{\left[\mathcal{L}\mathopen{}\mathclose{{\left(\mathbf{y},\;\Phi_{\theta,\omega_{j}}(\mathcal{P})}}\right)}}\right]$.
+PPT [82] instantiates $\{\omega_{j}\}_{j=1}^{|\mathbb{D}|}$ as the learnable gain and bias terms of the normalization layers of PTv3, trained on a per-dataset basis.
+During inference, the source dataset $\mathcal{D}_{j}$ of an incoming point cloud $\mathcal{P}$ must be identified in order to apply the correct normalization $\omega_{j}$, usually obtained from a lightweight dataset classifier as in One-for-All [75].
+Language-guided classification.
+To bridge label discrepancies across datasets (i.e.
+$\mathcal{C}_{i}\neq\mathcal{C}_{j}$ for datasets $\mathcal{D}_{i}$ and $\mathcal{D}_{j}$), we project per-point features to a shared language space using CLIP [57] text embeddings, following PPT [82].
+This enables supervision via class names, without relying on dataset labels.
+For instance, the class “pillow” is missing in ScanNet (grouped as “other”) but explicitly labeled in Structured3D; aligning to the CLIP embedding of “pillow” transfers knowledge across datasets even without direct supervision.
+All baselines and our Point-MoE are adapted to this language-guided protocol, ensuring fair comparison.
+3.2 Point-MoE
+Mixture-of-Experts (MoE) layer.
+Mixture-of-Experts [30] is a conditional computation architecture enabling significant capacity expansion while only modestly increasing computational cost.
 The core of MoE is to route input tokens dynamically to specialized subnetworks, referred to as experts, via a lightweight gating mechanism.
-Given an input feature vector $\mathbf{x}\in\mathbb{R}^{d}$, each MoE layer contains $N$ experts $\{f_{i}\}_{i=1}^{N}$ and a gating network $G:\mathbb{R}^{d}\rightarrow\mathbb{R}^{N}$.
+Given an input feature vector $\mathbf{x}\in\mathbb{R}^{D}$, each MoE layer contains $N$ experts $\{f_{i}\}_{i=1}^{N}$ where $f_{i}:\mathbb{R}^{D}\!\to\!\mathbb{R}^{D}$ and a gating network $G:\mathbb{R}^{D}\rightarrow\mathbb{R}^{N}$.
 For each input $\mathbf{x}$, the MoE layer output is computed as a weighted sum over the top-$k$ activated experts:
-$\mathrm{MoE}(\mathbf{x})=\sum_{i\in\mathcal{S}_{\mathbf{x}}}G_{i}(\mathbf{x})f_{i}(\mathbf{x})$ (1)
-where $\mathcal{S}_{\mathbf{x}}$ denotes the set of top-$k$ experts selected for input $\mathbf{x}$, and $G_{i}(\mathbf{x})$ represents the normalized gating weight assigned to expert $i$.
-In practice, the gating function is typically implemented via a linear projection followed by a sparse softmax, ensuring activation of only a small subset of experts, with $k\ll N$.
-Although auxiliary losses for balancing expert load are common in prior work [60], we empirically observed them to be unnecessary and therefore excluded them from our final setup.
+$\mathrm{MoE}(\mathbf{x})=\sum_{i\in\mathcal{S}_{\mathbf{x}}}G_{i}(\mathbf{x})\,f_{i}(\mathbf{x}),$ (1)
+where $\mathcal{S}_{\mathbf{x}}$ is the top-$k$ expert set and $G_{i}(\mathbf{x})$ are normalized gate weights.
+In practice, $G$ is a linear projection followed by a sparse softmax, activating only $k\!\ll\!N$ experts.
+Importantly, MoE introduces a large design space, including the choice of which components to replace, the number of experts, the routing strategy, whether to do auxiliary balancing loss and whether to use top-1 or top-2 gating.
+We systematically study how different MoE configurations affect performance on 3D point clouds through detailed ablations in the experiments.
 Integration into PTv3.
-To tailor the MoE architecture specifically for 3D point cloud data, we integrate the MoE mechanism directly into PTv3 by replacing the shared linear projection layers in PTv3’s attention module with an MoE layer.
-The original linear projections mapping input features into key, query, and value vectors within the attention mechanism are replaced by MoE layers, while all subsequent computations remain unchanged.
-This integration allows the MoE experts to specialize naturally in domain-specific transformations of point cloud data.
-Fig.
-1c provides a visual overview of the modified architecture.
-It is crucial to insert the MoE layer before the post-normalization; otherwise, we observe a significant reduction in performance gains.
-Language-guided classification.
-To bridge label discrepancies across datasets, we project per-point features to a shared language space using CLIP [56] text embeddings, following PPT [80].
-This enables supervision via class names, without relying on dataset labels.
-For instance, the class “pillow” is missing in ScanNet (grouped as “other”) but is explicitly labeled in Structured3D.
-Our model aligns point features with the CLIP [56] embedding of “pillow,” allowing knowledge to transfer across domains even in the absence of direct supervision.
-Domain-aware gating.
-Our preliminary experiments show that while basic Point-MoE naturally specializes experts implicitly, providing explicit domain information could further guide and accelerate expert specialization.
-To explore this, we introduce a dataset-aware gating variant termed Domain-Aware Point-MoE.
-In this variant, we explicitly associate each training domain $\mathcal{D}_{i}$ with a unique, learnable domain embedding vector $d_{i}\in\mathbb{R}^{d}$.
-Before routed through the gating mechanism, each input feature $x_{t}$ is concatenated with the corresponding domain embedding: $G_{t}(d)=\mathrm{softmax}(W\cdot(x_{t}\oplus d_{i}))$, where $W$ denotes a learnable linear projection, and the concatenation operation explicitly conditions gating decisions on the domain identity.
-This explicit conditioning encourages faster and more clearly delineated expert specialization aligned with dataset-specific characteristics observed during training.
-Domain randomization for robust gating.
-While domain-aware gating allows experts to specialize based on known domain identities, this approach assumes access to explicit domain labels at inference time.
-In practical scenarios, such labels are often unavailable—especially when deploying models in the wild, where domain boundaries may be ambiguous or entirely unknown.
-While a domain detector can be employed to approximate the domain label by mapping unseen domains to the closest known ones, such detectors may introduce errors or fail to capture nuanced domain shifts.
-This motivates the need for a more robust mechanism that does not rely on explicit domain identification at test time.
-To address this, we introduce a domain randomization strategy that enables the model to generalize to unseen or mixed-domain inputs without performance degradation on seen domains.
-Specifically, we define a learnable generic domain embedding and randomly assign it to 20% of the training points, replacing their true domain-specific embeddings.
-This forces the gating mechanism to learn domain-agnostic routing behavior, even in the presence of domain ambiguity.
-The remaining 80% of points retain their original domain embeddings, ensuring that the model still benefits from explicit supervision where available.
-In line with prior work [70, 17], we find that domain randomization improves generalization to unseen domains while preserving strong performance on those seen during training.
-Training objective.
-We use Mixture-of-Experts (MoE) in a supervised cross-domain setting, where the goal is to train a unified model $f(x;\theta)$ across $n$ domains $\{\mathcal{D}_{i}\}_{i=1}^{n}$, each containing samples $(X,Y)\sim\mathcal{D}_{i}$.
-The objective is to minimize the average loss across domains:
-$\min_{\theta}\sum_{i=1}^{n}\mathbb{E}_{(X,Y)\sim\mathcal{D}_{i}}\;\mathcal{L}(f(X;\theta),Y)$ (2)
-We consider two batching strategies during training: (1) Single-domain batching, where each batch is sampled from only one domain $\mathcal{D}_{i}$ at a time.
-(2) Mixed-domain batching, where each batch contains samples from multiple domains.
-Our implementation is based on the official open-source code of PTv3 released by Pointcept [20].
+Given a point cloud $\mathcal{P}$ with $M$ points, PTv3 [81] serializes and embeds the input into a token sequence $\mathbf{Z}=[z_{1},\ldots,z_{M}]$.
+PTv3 then applies stacked multi-head self-attention (MHSA) blocks interleaved with up- and down-sampling for multi-scale reasoning on these per-token embeddings.
+At any intermediate layer $\ell$, let $\mathbf{x}_{\ell}\!\in\!\mathbb{R}^{M\times D}$ denote the token sequence at layer $\ell$, where $D$ is the feature dimension, $H$ is the number of heads in multi-head attention, and $\mathrm{FFN}$ denotes a feed-forward network.
+The computations of a self-attention block using standard QKV-formulation and pre-norm residuals is as follows:
+$\mathbf{X}$ $=\mathrm{Norm}(\mathbf{x}_{\ell-1}),\qquad\mathbf{Q}=\mathbf{X}W_{q},\;\mathbf{K}=\mathbf{X}W_{k},\;\mathbf{V}=\mathbf{X}W_{v},$ (2)
+$\mathbf{A}$ $=\mathrm{Softmax}\!\big(\mathbf{Q}\mathbf{K}^{\top}/\sqrt{d_{h}}\big)\,\mathbf{V},\quad d_{h}=D/H,$ (3)
+$\mathbf{O}$ $=W_{o}\mathbf{A},$ (4)
+$\mathbf{x}^{\prime}_{\ell}$ $=\mathbf{x}_{\ell-1}+\mathbf{O},$ (5)
+$\mathbf{x}_{\ell}$ $=\mathbf{x}^{\prime}_{\ell}+\mathrm{FFN}\!\big(\mathrm{Norm}(\mathbf{x}^{\prime}_{\ell})\big)$ (6)
+The learnable query, key, and value projections are denoted as $W_{q}$, $W_{k}$ and $W_{v}$, while $W_{o}$ denotes the projection applied on the attention output $\mathbf{A}$.
+We apply MoE only at the attention output projection ($W_{o}$) in every block, rewriting Eq.
+(4) as $\mathbf{O}=\mathrm{MoE}(\mathbf{A})$ ; the query, key, and value projections ($W_{q},W_{k},W_{v}$) remain dense.
+Empirically, placing MoE at $W_{o}$ yields better performance than placing it in the FFN under the same expansion (Tab.
+3).
+We hypothesize that the gain comes from adding nonlinearity and capacity on the projection path: with purely linear projections, the composition of $W_{v}$ and $W_{o}$ can be merged into a single projection [47].
+Beyond MoE-layer placement, we employ mixed-dataset training: each minibatch jointly samples from both indoor and outdoor data (Tab.
+4).
+Having tokens from multiple datasets within a single minibatch, and thus contributing to the gradients of one update step, is useful in facilitating cross-sample interaction and emergent expert specialization.
+To illustrate the general applicability of this framework, we further extend MoE to OA-CNNs [52], a sparse convolution backbone.
+MoE-augmented OA-CNNs likewise outperform the dense version, further underscoring the effectiveness and general applicability of MoE as it is not tied to a specific backbone architecture.
+Detailed results are in the supplementary material under Sec.
+B.2.
 4 Experiments
-We conduct comprehensive experiments to evaluate Point-MoE for cross-domain 3D semantic segmentation.
+We conduct comprehensive experiments to evaluate Point-MoE for 3D semantic segmentation.
 First, we describe baseline methods and experimental setups (Sec.
@@
 4.1 Experimental Setup
-Baseline methods.
-Since our proposed Point-MoE framework is built upon PTv3 [78], we adopt PTv3 and PPT [80] as primary baselines.
-PPT, originally designed for dataset-specific BatchNorm  [34] and LayerNorm [4] layers and fine-tuning, is adapted here for joint training.
-Because PPT requires dataset labels at inference time, we follow One-for-All [74] to train a lightweight domain classifier (achieving over 95% accuracy), which we also use for Point-MoE under the domain-aware setting.
-To ensure a fair comparison, we begin by benchmarking the dense PTv3 model under the domain-agnostic joint training setting, systematically varying key design factors such as mixed-batch composition and normalization strategy.
-In this setup, models are trained across multiple domains without access to dataset labels, and no label information is provided at inference time.
-This setting reflects a realistic deployment scenario where domain metadata is unavailable.
-For the domain-aware joint training setting, we compare to the PPT + One-for-All baseline.
-Here, models are trained with access to dataset labels.
-During inference, we either provide predicted labels using a trained domain classifier or assign a “generic” class to simulate deployment in unknown or mixed domains.
-This allows us to isolate the value of domain conditioning and assess the robustness of Point-MoE under both domain-aware and domain-agnostic regimes.
 Datasets.
-Experiments are conducted on three indoor datasets—ScanNet [21], S3DIS [2], and Structured3D [88]—and evaluated on both seen and unseen domains, including Matterport3D [12].
-We then evaluate the best-performing Point-MoE and baseline configurations under a more challenging indoor-outdoor setting, incorporating outdoor datasets from nuScenes [11], SemanticKITTI [7], and Waymo [65].
-Voxel sizes are standardized to 0.02m for indoor scenes and 0.05m for outdoor scenes.
-Throughout, we denote model capacity using -S (small) and -L (large) suffixes for all architectures, including PTv3, Point-MoE, and PPT.
+We conduct experiments on three indoor benchmarks: ScanNet [19], S3DIS [1], and Structured3D [95], and additionally evaluate zero-shot on Matterport3D [10].
+Specifically, we train a single model on ScanNet, S3DIS, and Structured3D, and directly evaluate its performance on both seen and unseen datasets via the language-guided head using the same checkpoint.
+We further test the best-performing Point-MoE architectures and baseline configurations in a more challenging indoor and outdoor joint setting that incorporates nuScenes [9] and SemanticKITTI [5] as training data, and Waymo [67] for zero-shot evaluation.
+The same evaluation protocol is used across both settings.
+Baselines.
+Since Point-MoE is built on PTv3 [81], we adopt PTv3 and PPT [82] as our primary baselines.
+PPT, originally designed with dataset-specific normalization and per-dataset fine-tuning, is adapted here to the joint training setting.
+Because PPT requires dataset labels at inference, we follow One-for-All [75] and train a lightweight dataset classifier (achieving $\geq$95% accuracy) to predict the source dataset for an input sample at test time (Sec.
+B.2).
+Evaluation and Implementation Details.
+To ensure fair comparison, we re-benchmark PTv3 under our unified protocol by varying batch composition and normalization, which yields an average improvement of 22 mIoU (Sec.
+B.3).
+We also compare against PPT combined with One-for-All, which uses ground-truth dataset labels during training and predicted dataset labels during inference.
+Model capacity is denoted with -S (small) and -L (large) suffixes for all architectures, including PTv3, Point-MoE, and PPT.
+PTv3-S and PPT-S use 14 encoder and 8 decoder layers, while their -L variants use 18 encoder and 12 decoder layers; in contrast, both Point-MoE-S and -L have 14 encoder and 8 decoder layers, with -S using 4 experts per MoE layer and -L using 8.
+The expert hidden dimension is set to 1× the input dimension in Point-MoE-S and 2× in Point-MoE-L.
+For details, please refer to Tab.
+10 in the supplementary material.
 4.2 Experimental Results
-Table 1: Performance of Joint Training for Indoor 3D Semantic Segmentation.
-All models are jointly trained on three indoor datasets and evaluated under two settings.
-In Domain-agnostic joint training, models are evaluated using only the point cloud and the corresponding dataset-specific label set.
-In Domain-aware joint training, a predicted scene label from a domain classifier is provided at inference time to guide the model.
-“-G” indicates that model utilizes domain randomization and contains a learnable “generic” domain label during training.
-All results are reported as mean Intersection-over-Union (mIoU).
-Methods Params Activated Mix Layer. ScanNet Structured3D S3DIS Matterport3D
-Seen Unseen
-Val Val Test Val Val Test
+Table 1: Results on seen datasets.
+Results are reported in mIoU.
+Point-MoE achieves the best average performance without using domain labels.
+Methods marked with use dataset labels during training.
+Models trained on a single dataset but evaluated on unseen datasets are in gray.
+(a) Indoor-only joint training on ScanNet [19], Structured3D [95], and S3DIS [1].
+(b) Joint training across indoor and outdoor datasets: nuScenes [9] and SemanticKITTI [5].
+[…]
+(a) Indoor-only joint training
+Methods Params Activated ScanNet Structured3D S3DIS Avg.
+Val Val Test Area5
 Single-dataset Training
-PTv3-ScanNet [78] 46M 46M 75.0 19.8 20.9 8.8 38.9 44.5
-PTv3-Structured3D [78] 46M 46M 21.2 81.5 79.4 7.1 21.0 24.0
-PTv3-S3DIS [78] 46M 46M 4.9 3.5 5.3 67.6 5.0 3.4
-PTv3-Matterport3D [78] 46M 46M 53.1 22.9 23.7 9.8 54.1 56.0
-Domain-agnostic joint training
+PTv3-ScanNet [81] 46M 46M 75.0 19.8 20.9 8.8 31.1
+PTv3-Structured3D [81] 46M 46M 21.2 81.5 79.4 7.1 47.3
+PTv3-S3DIS [81] 46M 46M 4.9 3.5 5.3 67.6 20.3
+Multi-dataset Joint Training
+OA-CNNs [52] 52M 52M 71.6 59.0 58.6 60.6 62.5
+PTv3-S [81] 46M 46M 71.3 56.6 55.9 69.4 63.3
+PTv3-L [81] 97M 97M 72.2 56.0 58.1 67.4 63.4
+PPT-S [82, 75] 47M 47M 74.7 64.2 65.5 64.7 67.3
+PPT-L [82, 75] 98M 98M 74.8 64.7 65.9 65.0 67.6
+Point-MoE-S 59M 52M 74.5 69.0 69.9 68.0 70.4
+Point-MoE-L 100M 59M 76.0 69.5 70.8 69.5 71.5
 […]
+(b) Indoor and Outdoor joint training
+Methods Params Activated ScanNet Structured3D S3DIS nuScenes SemKITTI Avg.
+Val Val Test Area5 Val Val
+PTv3-L [81] 97M 97M 74.6 64.1 63.9 68.0 69.6 62.9 67.2
+PPT-L [75, 82] 98M 98M 76.7 67.5 66.0 61.9 70.8 66.9 68.3
+Point-MoE-L 100M 59M 77.4 69.5 70.2 68.1 72.9 66.4 70.8
+Table 2: Zero-shot results on unseen datasets.
+Point-MoE consistently achieves the strongest generalization across unseen datasets in the zero-shot setting.
+Methods marked with use domain labels during training.
 […]
+(a) Indoor-only joint training
+Method Matterport3D Average
+Val Test
+PTv3-L 39.0 45.4 42.2
+PPT-L 27.8 26.3 27.1
+Point-MoE-L 41.0 46.7 43.9
 […]
-[…]
-[…]
-[…]
-[…]
-[…]
-Domain-aware joint training
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-Table 2: Cross-domain 3D Semantic Segmentation across Indoor and Outdoor Scenes.
-We evaluate the best-performing models from the indoor-only setting in a more challenging joint training setup that spans both indoor and outdoor datasets.
-This setup better reflects real-world deployment scenarios, requiring models to generalize across diverse domains.
-ScanNet Structured3D S3DIS nuScenes SemKITTI Matterport3D Waymo
-Indoor Outdoor Indoor Outdoor
-Seen Unseen
-Methods Params Activated Mix Layer. Val Val Test Val Val Val Val Test Val Test
-Domain-agnostic joint training
-[…]
-[…]
-[…]
-[…]
-Domain-aware joint training
-[…]
-[…]
-[…]
-Single-dataset training.
-We begin by evaluating the standard setting in which models are trained and evaluated on individual datasets.
-Note we train end evaluate PTv3 with language-guided classification.
+(b) Indoor and Outdoor joint training
+Method Matterport3D Waymo Average
+Val Test Val Test
+PTv3-L 39.1 45.9 21.9 23.2 32.5
+PPT-L 24.9 23.1 16.3 16.7 20.3
+Point-MoE-L 41.3 49.6 23.7 25.3 35.0
+Single-dataset training generalizes poorly.
+We begin by evaluating the standard setting where models are trained and evaluated on individual datasets.
+Note that in our setup PTv3 is trained and evaluated with CLIP-guided classification in an open-set setting, consistent with the multi-dataset regime [82], whereas the original PTv3 paper reports results under standard closed-set softmax label classification [81].
+This difference makes our scores not directly comparable to the numbers in the original PTv3 paper.
 As shown in Tab.
-1, PTv3 achieves strong in-domain performance, but its performance drops significantly on unseen domains.
+11, PTv3 achieves strong results on seen datasets, but its performance drops significantly on zero-shot datasets.
 For instance, a PTv3 model trained on S3DIS achieves only 3.5% on Structured3D and 4.9% on ScanNet.
-While domain-specific models are not expected to generalize across domains or categories, these results underscore the brittleness of single-dataset training.
-This motivates the need for cross-domain training strategies that expose the model to diverse data and promote broader generalization.
-Domain-agnostic joint training.
-In the indoor-only setting, Point-MoE outperforms all variants of PTv3 in most domains and performs competitively on S3DIS.
-Referring to Tab.
-1, Point-MoE-S achieves 75.4% on ScanNet and 69.7% on S3DIS, outperforming PTv3-S by +4.1% and +13.8%, respectively.
-Remarkably, Point-MoE-S activates only 52M parameters compared to 97M in PTv3-L, yet still surpasses PTv3-L by +0.8% on ScanNet and +8.4% on S3DIS.
-In addition, Point-MoE-L reaches 75.8% on ScanNet and 73.0% on Structured3D, significantly ahead of PTv3-L.
-On the unseen Matterport3D domain, Point-MoE-L obtains 41.6% on validation set and 47.8% on test set, outperforming PTv3-L by +1.4% and +1.2% respectively while activating only 59M parameters at inference.
-In the more challenging indoor and outdoor setting, Point-MoE-L again leads across the board with a few being competitve.
-Referring Tab.
-2, Point-MoE-L achieves 69.2% on S3DIS validation and 70.2% on test split, outperforming the best dense baseline by +5.4% and +6.3%, respectively.
-On outdoor datasets, Point-MoE-L achieves 72.9% on nuScenes exceeding the dense baseline by +2.6%.
-Notably, on the unseen datasets, Point-MoE-L outperforming all dense PTv3 variants by a clear margin.
-These gains are achieved while activating nearly half the parameters of PTv3-L, underscoring the efficiency and scalability of sparse expert routing.
-Domain-aware joint training.
-We further evaluate Point-MoE in the domain-aware setting, where models are trained with access to dataset labels.
-In the indoor-only setting, both domain-aware Point-MoE and its generic variant (Point-MoE-G) surpass all strong baselines: variants of PPT [80] + One-for-All [74].
+While dataset-specific models are not expected to generalize across datasets or categories, these results underscore the brittleness of single-dataset training.
+This motivates the need for multi-dataset joint training strategies that expose the model to diverse data and promote broader generalization.
+Point-MoE surpasses all baselines on seen datasets, without dataset labels during training.
+Point-MoE dispenses with dataset identifiers during training yet yields consistent gains across indoor benchmarks (Tab.
+11); both Point-MoE-S and Point-MoE-L variants improve on ScanNet, Structured3D, and S3DIS, with Point-MoE-L attaining a new state-of-the-art average of $71.5$ mIoU.
+Despite joint training over multiple datasets, Point-MoE matches or exceeds single-dataset specialists, indicating effective expert routing rather than overfitting to any one source, or significantly exhibiting the negative transfer effects noted in PPT [82].
+In the more challenging indoor and outdoor joint training setting (Tab.
+11), Point-MoE maintains its advantage across datasets and, on average, surpasses PTv3-L by $3.55$ mIoU and PPT-L by $2.45$ mIoU, demonstrating that label-free expert routing scales to heterogeneous datasets while simplifying the training pipeline.
+Moreover, performance on the seen indoor datasets remains closely aligned between indoor-only training and indoor-outdoor joint training, suggesting that Point-MoE preserves seen datasets accuracy even when subsequently trained jointly with outdoor data.
+This generalization is achieved implicitly in the model, without requiring dataset labels for incoming samples or maintaining explicit per-dataset normalization parameters.
+Note, since we use a lightweight dataset classifier, PPT’s performance differs from what is reported in the original paper, where oracle dataset identifiers are provided at inference.
+Finally, because only a subset of experts is activated, Point-MoE achieves a favorable accuracy–compute trade-off, cutting compute by $30.9\%$ (265.7 vs.
+384.4 GFLOPs) and peak VRAM by $19.0\%$ (33.3 vs.
+41.1 GiB) relative to PPT-L (Tab.
+5).
+We analyze the emergent dataset-awareness and routing behavior in Sec.
+4.
+Point-MoE outperforms all baselines consistently across zero-shot datasets.
 As shown in Tab.
-1, Point-MoE-G-L achieves 76.7% on ScanNet, 68.0% on Structured3D, and 67.0% on S3DIS, outperforming the best PPT + One-for-All baseline by +2.0%, +3.1%, and +2.3%, respectively.
-On the unseen Matterport3D dataset, it reaches 41.8% on the validation split and 47.8% on the test set, again outperforming all baselines.
-Notably, using only 20% of training data tagged as “generic” does not degrade performance on seen domains.
-In the indoor-and-outdoor setting, Point-MoE-G-L continues to outperform all baselines across the board.
-Referring to Tab.
-2, it achieves 77.0% on ScanNet, 72.1% and 74.2% on Structured3D (val/test), 62.4% on S3DIS and 72.7% on nuScenes—surpassing the best baseline by +0.3%, +4.6%, +4.1%, +1.0% and +0.5%, respectively.
-On the unseen domains, including Matterport3D and Waymo, Point-MoE-G-L delivers the strongest generalization.
-Crucially, it performs on par with domain-aware Point-MoE on seen datasets while achieving superior accuracy on unseen ones.
-These results validate that domain randomization with a generic label can enhance robustness without compromising in-domain performance.
-We still observe a performance gap between cross-domain and single-dataset training in cases like Structured3D [88].
-Single-dataset models can adapt to domain-specific cues and distributions, while Point-MoE must balance capacity across domains.
-As a result, its performance may trail on large, homogeneous datasets, but remains competitive overall due to its generalization ability.
-4.3 MoE Design Ablation
-Table 3: Ablation of Point-MoE Design and Hyperparameters.
+22 and Tab.
+22, Point-MoE-L achieves the strongest zero shot performance, including on Matterport3D under indoor only training and on both Matterport3D and Waymo under indoor and outdoor training, without using dataset labels during training.
+We believe a key reason is that label free expert routing encourages specialization based on the underlying semantics and geometry of the input rather than on dataset identity, resulting in more stable behavior under distribution shift.
+By contrast, PPT-L relies on explicit dataset labels during training.
+While this provides useful supervision on seen datasets, it can also encourage the model to depend too heavily on dataset specific cues.
+As a result, when a dataset is held out and its label is unavailable or no longer aligned with the training distribution at test time, this dependence becomes brittle and can hurt generalization.
+Additional analysis of Point-MoE’s zero shot adaptation is provided in Sec.
+4.4.
+Table 3: Ablations of Point-MoE designs and hyperparameters.
 We conduct a systematic ablation on key design and hyperparameter choices of Point-MoE across four indoor multi-domain datasets: ScanNet (Scan), Structured3D (S3D), S3DIS, and Matterport3D (Mat).
-[…]
-[…]
+$\alpha$ Scan S3D S3DIS Mat
+0 74.5 66.9 68.6 40.1
 $10^{-4}$ 72.3 64.1 67.9 39.5
@@
 (a) Auxiliary load balancing loss. Removing the auxiliary loss consistently improves performance.
-[…]
+top $k$ Scan S3D S3DIS Mat
 top 1 74.4 64.8 65.5 39.0
-[…]
+top 2 74.5 66.9 68.6 40.1
 top 3 73.8 62.4 67.8 39.7
-(b) Top-$k$ expert selection. Best performance is achieved with $k{=}2$ experts.
-[…]
-[…]
+(b) Top-$k$ expert selection. Best performance is achieved with k=2k{=}2 experts across datasets.
+norm Scan S3D S3DIS Mat
+BN. 74.5 66.9 68.6 40.1
 LN. 70.8 56.9 67.3 37.2
@@
 (c) Normalization. BatchNorm yields strongest performance; RMSNorm degrades results.
-[…]
-dec. 74.0 64.9 66.6 38.1
-[…]
-(d) MoE position. Inserting MoE in all layers yields the best results.
-[…]
-[…]
-[…]
+case Scan S3D S3DIS Mat
+FFN 72.6 66.1 66.9 39.2
+Proj. 74.5 66.9 68.6 40.1
+(d) MoE position. Replacing the projection layer performs better.
+act. Scan S3D S3DIS Mat
+SiLU 73.6 65.7 64.9 40.1
+ReLU 74.5 66.9 68.6 40.1
 (e) Activation functions. ReLU performs the best.
-[…]
-[…]
-[…]
+case Scan S3D S3DIS Mat
+1H 73.0 60.3 66.3 39.5
+2H 73.5 58.6 65.4 39.8
 (f) Expert width. H indicates original feature dimension.
-[…]
-[…]
+shared Scan S3D S3DIS Mat
+0 74.5 66.9 68.6 40.1
 1 73.6 62.3 66.6 39.1
 (g) Shared experts. Not sharing experts yields better performance.
-[…]
-[…]
-[…]
+num. Scan S3D S3DIS Mat
+4 74.5 66.9 68.5 40.1
+8 73.6 68.7 68.5 40.4
 (h) Number of experts. More experts improve performance.
-[…]
+batch Scan S3D S3DIS Mat
 4 74.0 64.9 66.6 38.1
-[…]
+6 74.3 66.9 68.6 40.1
 (i) Batch size. Larger batchsize improves performance.
-We investigate the impact of MoE design choices and hyper-parameters choices in indoor cross-domain 3D semantic segmentation.
-All experiments are conducted with ScanNet [21], Structured3D [88] and S3DIS [2] and evaluate without providing dataset label on seen dataset and also evaluate on unseen dataset: Matterport3D [12].
+4.3 MoE Design Ablation
+We investigate the impact of MoE design choices and hyper-parameters choices in indoor multi-dataset 3D semantic segmentation.
 Results are presented in Tab.
 3.
-Auxiliary load-balancing loss.
+Note we are not using precise evaluator so results are not directly comparable to Tab.
+1.
+Please refer to details in Sec.
+B.3 in the supplementary material.
+Effect of load-balancing loss.
 The auxiliary load-balancing loss is designed to encourage uniform expert usage and prevent expert collapse [56, 78, 42].
@@
 3.
-Interestingly, we find that removing the auxiliary loss entirely leads to consistently better performance across domains.
+Interestingly, we find that removing the auxiliary loss entirely leads to consistently better performance across datasets.
 In contrast, stronger auxiliary losses result in significant degradation.
-We hypothesize that this is due to the inherent imbalance in the distribution of samples across domains in 3D point cloud datasets—a phenomenon also observed by ChartMoE [82].
+We hypothesize that this is due to the inherent imbalance in the distribution of samples across datasets in 3D point cloud datasets—a phenomenon also observed by ChartMoE [87].
 Effect of activated experts.
@@
 3 compares the performance of configurations with 1, 2 and 3 activated experts, finding diminishing return when increasing number of activated experts.
-We find that activating two experts yields the best performance across domains.
+We find that activating two experts yields the best performance across datasets.
 This analysis suggests that merely increasing the number of activated experts does not guarantee improved performance.
-MoE position.
-We ablate the placement of MoE layers to understand where expert specialization is most effective.
-Empirically, we observe greater expert diversity when MoE is applied in the decoder compared to the encoder.
-In Tab.
-3, we compare two configurations: applying MoE layers only in the decoder, versus applying them in both the encoder and decoder.
-Results show that applying MoE to both stages consistently yields better performance, suggesting that expert specialization benefits from exposure to both encoder and decoder.
-Effect of scaling model with MoE.
+Table 4: Ablations for mixed-dataset training. “Mix” indicates that each training batch combines samples from multiple datasets.
+Methods Mix Scan S3D S3DIS Mat Average
+PTv3 48.4 57.8 44.1 27.8 44.5
+[…]
+PPT 74.2 58.2 61.0 24.3 54.4
+[…]
+PT-MoE 54.1 60.9 40.6 28.0 45.9
+[…]
+Effect of MoE position.
+We compare placing MoE in the feed-forward network (FFN) versus in the attention output projection $W_{o}$, holding all other components fixed.
+Projection-MoE consistently outperforms FFN-MoE in Tab.
+3.
+We hypothesize that $W_{o}$ is the fusion point where multi-head outputs are recombined before the next normalization, so experts placed there can route on richer, head-aggregated signals that preserve cross-dataset geometry cues, while leaving $W_{q},W_{k},W_{v}$ shared and stable; in contrast, FFN experts operate on post-normalized activations where such dataset or geometry cues are more attenuated.
+This empirical gain leads us to adopt projection-MoE as the default.
+Similar observations also appear in recent work: e.g.
+UMoE [89] finds that using MoE for the output (and value) projections yields strong compute savings with high performance.
+Effect of model scaling with MoE.
 Tab.
 3 compares Point-MoE with 4 and 8 experts with activated experts fixed at 2.
-Increasing the number of experts from 4 to 8 leads to improved performance on seen domains, and yields a modest gain of 0.4% on the unseen domain.
+Increasing the number of experts from 4 to 8 leads to improved performance on seen and zero-shot datasets.
 Tab.
-3 ablates the expert width by varying the intermediate size of the MLP (e.g., 1H, 2H).
-The results show a mixed trend: wider experts improve performance on ScanNet and Matterport3D, but degrade results on Structured3D and S3DIS.
-Finally, we investigate the effect of adding shared experts [63, 72, 22], which are applied universally without gating.
-We find that disabling shared experts consistently leads to better accuracy and domain specialization across all benchmarks, as reported in Tab.
+3 ablates the expert width by varying the intermediate feature dimension and the results show a mixed trend: wider experts improve performance on ScanNet and Matterport3D, but degrade results on Structured3D and S3DIS.
+Finally, we investigate the effect of adding shared experts [65, 73, 20].
+We find that disabling shared experts consistently leads to better accuracy and dataset specialization across all benchmarks, as reported in Tab.
 3.
-Ablation of hyper-parameters choices.
-We provide justification for our selected hyperparameters.
+Effect of hyper-parameter choices.
+We provide justification for our selected hyper-parameters.
 In Tab.
-3, BatchNorm [34] consistently outperforms other normalization methods [4, 85] in the cross-domain setting.
+3, BatchNorm [29] consistently outperforms other normalization methods [3, 91] in the multi-datasets setting.
 Tab.
@@
 3 demonstrates that importance of batch size in MoE training.
-Specifically, the batch size must exceed the number of experts to avoid performance degradation.
-In general, larger batch sizes lead to improved performance across domains.
+In general, larger batch sizes lead to improved performance across datasets.
+Mixed dataset training.
+In joint multi-dataset training, forming each minibatch with samples drawn from multiple datasets is consistently beneficial.
+Although there is no explicit batch-level interaction module, samples from different datasets still influence one another through shared normalization statistics and, for Point-MoE, through competition over a common expert pool.
+As shown in Tab.
+4, this mixed-dataset training strategy improves the average mIoU of all backbones: PTv3 +8.3+8.3, PPT +2.9+2.9, and Point-MoE +16.6+16.6 mIoU on average.
+For dense backbones such as PTv3 and PPT, mixing exposes the model to more diverse domain statistics within each optimization step, which appears to regularize training and reduce overfitting to dataset-specific feature distributions.
+To make this comparison fair for PPT, we re-implemented its normalization module so that mixed-dataset batching is supported.
+The effect is most pronounced for Point-MoE, where diverse minibatches promote expert specialization across datasets, stabilizing routing and improving generalization.
+Figure 2: t-SNE visualization of feature clustering.
+The first three columns show that the decoders of both PPT and PointMoE have better separation between datasets in their decoder representations, than vanilla PTv3.
+The rightmost column highlights how Point-MoE generalizes to zero-shot datasets, with Matterport3D features aligning closely with ScanNet, which is indeed semantically similar.
 4.4 Analysis of MoE
-Figure 2: Training Loss and Validation mIoU over time.
-Left: training loss curves for joint training across multiple domains.
-Right: validation mIoU on indoor datasets.
-Point-MoE-L achieves competitive performance to PPT, and both significantly outperform the baseline PTv3-L.
-Training efficiency and validation mIoU.
+Dataset clustering emerges in Point-MoE.
 Fig.
-2 shows the training loss and validation mIoU curves for four models: the baseline PTv3-L, its improved variant PTv3-Mix-LN (which incorporates mixed-domain batches and LayerNorm), PPT-L [80], and our proposed Point-MoE-L.
-All models are trained from scratch across multiple domains.
-Point-MoE-L converges faster and achieves strong validation mIoU without using explicit dataset labels, matching or exceeding the performance of PPT-L trained with ground-truth domain labels.
-While all models reach similar training loss, only Point-MoE-L and PPT-L generalize effectively—reinforcing that low training loss is not indicative of strong cross-domain performance.
-On ScanNet, Structured3D, Matterport3D, and nuScenes, Point-MoE-L shows consistent improvement and avoids early plateaus seen in PPT-L, especially on Structured3D, suggesting stronger long-term learning.
-PTv3-L fails to generalize and exhibits unstable validation curves.
-PTv3-Mix-LN improves stability over PTv3-L significantly, but still underperforms Point-MoE-L and PPT-L.
-Figure 3: Expert Choices Visualization.
-Different color represents a different expert.
-(a) shows a specific expert focusing on edges across the scene, revealing spatially coherent routing.
-(b) and (c) highlight expert selections in indoor scenes, where semantically related regions (e.g., chair, desk) are consistently routed to the same expert.
-(d) displays an outdoor scene with sparse LiDAR points, yet the model still demonstrates meaningful routing despite limited geometry.
+2 shows t-SNE visualizations [70] of encoder and decoder features from PTv3, PPT [82], and Point-MoE.
+PTv3 exhibits no clear separation in either space, indicating limited capacity to form dataset-specific representations.
+PPT, by contrast, produces sharply separated clusters at both encoder and decoder levels, reflecting its reliance on dataset labels during training.
+While this aids seen-datasets performance, such rigid partitioning may hinder generalization capabilities as highlighted in Tab.
+2.
+Point-MoE displays a different pattern: encoder features remain mixed across datasets, supporting shared representation learning, while the decoder disentangles dataset-specific structures.
+This apparent division of labor suggests that Point-MoE performs implicit dataset inference, without requiring explicit labels.
+Crucially, on the zero-shot Matterport3D dataset, Point-MoE aligns samples with the most related seen-dataset clusters (e.g., ScanNet), which likely contributes to its strong generalization.
+Overall, these results show that Point-MoE can self-organize its experts and adapt to novel environments by effectively transferring knowledge across diverse training datasets.
+Figure 3: Expert choice visualization.
+Different colors within each image indicate different experts, showing that Point-MoE self-organizes point routing based on spatial and semantic cues across different layers.
+(a) shows one expert (green) focusing on edges, suggesting spatially aware mid-level routing in the encoder.
+(b) and (c) show that semantically related regions (e.g., chairs and desks) are consistently assigned to the same expert, even across datasets.
+(d) shows meaningful routing in an outdoor scene with sparse LiDAR points despite lower point density and more irregular geometry.
 Expert choice visualization.
 Fig.
-3 showcases expert assignments for a validation scene at selected layers.
-In (a), we observe that early encoder layers rely heavily on geometric cues for routing.
-For instance, green experts are consistently activated along object boundaries such as the edges of desks and chairs, while red experts dominate flat surfaces.
-In (b) and (c), the decoder layers exhibit more semantically meaningful expert selection—likely due to their proximity to the loss function—with distinct experts attending to objects like desks, chairs, floors, and walls.
-In (d), we examine an outdoor scene with sparse LiDAR data.
-Despite limited geometric structure, the model still organizes routing meaningfully: nearby points are routed to blue experts, while farther points activate red experts.
-We include more visualizations in the appendix for completeness.
-We also note occasional visual artifacts where isolated points are assigned different experts than their neighbors, which may be related to PTv3’s architectural choices such as point serialization or positional encoding.
-Figure 4: Expert routing behavior across domains.
-The left panel shows the most frequently used expert paths through each encoder (E-) and decoder (D-) layer, with feature channel sizes marked in parentheses, illustrating the variation in expert selection for tokens from different datasets.
-The router tends to assign experts based on the input domain, particularly in the decoder MoE layers.
-To quantify this, we compute the Jensen-Shannon Divergence (JSD) between the expert selection distributions of different datasets at each MoE layer.
+3 visualizes expert assignments for validation scenes at selected encoder and decoder layers.
+In (a), the encoder layer primarily routes points according to local geometric structure.
+In particular, one expert (green) is consistently activated around object boundaries and thin structural regions, such as the edges of desks and chairs, while other experts dominate flatter surfaces, suggesting spatially aware mid-level routing in the encoder.
+In (b) and (c), the decoder layers exhibit more semantically coherent expert selection.
+Semantically related regions, such as chairs, desks, floors, and walls, are more consistently assigned to the same experts, even across different scenes and datasets, likely because decoder features are closer to the supervision signal and thus encode higher-level semantics.
+In (d), we examine an outdoor scene with sparse LiDAR points.
+Despite the lower point density and more irregular geometry, the model still produces meaningful routing patterns rather than random assignments.
+Nearby foreground regions and farther background regions tend to activate different experts, suggesting that the learned routing generalizes beyond dense indoor scans to more challenging outdoor environments.
+Additional visualizations are provided in Sec.
+C.3.
+We also observe a few isolated points assigned to different experts than their neighbors, possibly due to PTv3 design choices such as point serialization and positional encoding.
+Table 5: Inference efficiency. FLOPs per step are measured with a profiler, and peak VRAM reports allocated memory. Relative deltas are computed with respect to PPT-L.
+Model FLOPs/step (GFLOPs) Peak VRAM (GiB) $\Delta$ FLOPs vs. PPT-L $\Delta$ VRAM vs. PPT-L
+PPT-L 384.4 41.1 0% 0%
+PTv3 375.8 37.4  $\downarrow\,2.2\%$  $\downarrow\,9.0\%$
+Point-MoE-L 265.7 33.3  $\downarrow\,30.9\%$  $\downarrow\,19.0\%$
+Figure 4: Expert routing across datasets.
+Left: the most frequent expert paths through each encoder (E) and decoder (D) layer, with channel sizes in parentheses, showing that expert selection varies across datasets.
+The effect is especially pronounced in decoder MoE layers.
+Right: Jensen Shannon Divergence (JSD) between dataset specific expert selection distributions at each MoE layer.
 Token pathways.
-To understand how Point-MoE adapts to diverse domains, we analyze expert routing behavior at the token level.
-Specifically, we track the top-1 expert assignment for each token across all MoE layers and construct full routing trajectories from the final layer back to the first.
-We then identify the top 100 most frequent expert paths based on their occurrence across all tokens.
-As shown in Fig.
-4, encoder expert paths are substantially less diverse than those in the decoder, indicating that encoder layers perform more domain-agnostic processing.
-Interestingly, we observe a sparse routing pattern in deeper encoder layers.
-This may be attributed not to feature reuse, but to the U-Net-style design where features are spatially deep but token sparsity increases, resulting in reduced variability in routing decisions.
-When examining domain-level trends, we find that certain dataset pairs—such as SemanticKITTI and nuScenes or ScanNet and Structured3D—share similar expert pathways, suggesting that Point-MoE implicitly clusters domains with related geometric or semantic structures.
+To understand how Point-MoE adapts to diverse datasets, we analyze expert routing at the token level.
+We track each token’s top-1 expert across all MoE layers and reconstruct full routing trajectories from the final layer back to the first.
+We then rank expert paths by frequency and focus on the top 100 (Fig.
+4).
+Encoder paths are notably less diverse than decoder paths, suggesting that encoders perform more dataset-agnostic processing.
+We also find sparse routing in deeper encoder layers, likely due not to feature reuse but to the U-Net design: features grow deeper spatially while tokens become sparser, reducing routing variability.
+When examining dataset-level trends, we find that certain dataset pairs, such as SemanticKITTI and nuScenes or ScanNet and Structured3D, share similar expert pathways, suggesting that Point-MoE implicitly clusters datasets with related geometric or semantic structures.
 To quantify these observations, we compute the Jensen-Shannon Divergence (JSD) [45] between expert selection distributions across datasets at each MoE layer.
@@
 As shown in Fig.
-4 (right), decoder layers exhibit significantly higher JSD, indicating stronger domain-specific specialization.
-Several encoder layers also display nontrivial JSD, underscoring the benefit of placing MoE throughout the network.
+4 (right), decoder layers exhibit significantly higher JSD, indicating stronger dataset-specific specialization.
+Several encoder layers also display nontrivial JSD, underscoring the benefit of placing MoE throughout the network (see Sec.
+C.2 for more analysis).
+Efficiency analysis.
+As shown in Tab.
+5, Point-MoE-L requires only 265.7 GFLOPs per step and 33.3 GiB peak VRAM, reducing compute by $30.9\%$ and memory by $19.0\%$ relative to PPT-L (384.4 GFLOPs and 41.1 GiB).
+It is also more efficient than PTv3 (375.8 GFLOPs and 37.4 GiB).
+Despite this lower cost, Point-MoE-L achieves the best average mIoU in both settings, reaching $71.5$ in indoor only joint training and $70.2$ in indoor and outdoor joint training (Tab.
+1).
 5 Discussions
-This work embraces the “bitter lesson” [66] in AI: scalable generalization emerges from flexible architectures trained on diverse data, rather than hand-engineered domain priors [68].
+This work embraces the “bitter lesson” [68] in AI: scalable generalization emerges from flexible architectures trained on diverse data, rather than hand-engineered domain priors.
 We introduce Point-MoE, a mixture-of-experts architecture for 3D understanding that dynamically routes tokens to specialized experts across both indoor and outdoor domains.
-Our experiments demonstrate that Point-MoE not only achieves competitive performance on seen datasets but, crucially, also generalizes effectively to novel domains without explicit domain labels or supervision.
+Our experiments demonstrate that Point-MoE not only achieves superior performance on seen datasets but also generalizes effectively to zero-shot datasets without explicit domain labels.
 Through careful analysis of expert routing behavior, we show that the model self-organizes around both geometric and semantic structures, adapting its computation pathways to the nature of the input.
 These results highlight the potential of sparse, modular computation for robust and scalable 3D scene understanding.
-We provide more analysis of Point-MoE and discuss our limitations in the supplementary material.
+We provide more analysis of Point-MoE and discuss our limitations in the supplementary materials.
 6 Acknowledgement
-The authors gratefully acknowledge Research Computing at the University of Virginia for providing the computational resources and technical support that made the results in this work possible.
-Zezhou Cheng acknowledges the Adobe Research Gift, as well as GPU support provided by the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program and the National Artificial Intelligence Research Resource (NAIRR) Pilot program.
+The authors acknowledge the MathWorks Research Award, Adobe Research Gift, the University of Virginia Research Computing and Data Analytics Center, Advanced Micro Devices AI and HPC Cluster Program, Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, and National Artificial Intelligence Research Resource (NAIRR) Pilot for computational resources, including the Anvil supercomputer (National Science Foundation award OAC 2005632) at Purdue University and the Delta and DeltaAI advanced computing resources (National Science Foundation award OAC 2005572).
 Appendix
 For a thorough understanding of our Point-MoE, we have compiled a detailed Appendix. The table of contents below offers a quick overview and will guide to specific sections of interest.
-[…]
-[…]
-[…]
-[…]
 Appendix A Limitations, Licenses and Risks
 A.1 Limitations
-Our work is limited by computational resource constraints, which restrict the scale of both the Point-MoE architecture and the training datasets.
-While our results already demonstrate strong generalization across diverse 3D domains, larger models and a broader training corpus could further enhance performance.
-Additionally, our expert routing relies solely on the standard gating loss without auxiliary regularization for load balancing or diversity, which may lead to suboptimal expert specialization.
-Finally, the dynamic routing mechanism incurs nontrivial communication and synchronization overhead compared to dense backbones like Point Transformer V3, highlighting the need for more efficient expert parallelism or conditional computation in future work.
+Within our compute budget, we present a comprehensive study of MoE for point clouds—covering placement, expansion, routing choices (top-$k$, capacity factor, load balancing), and mixed-domain batching—with consistent gains.
+However, compute remains the primary bottleneck, capping Point-MoE’s scale (experts, width, depth), training schedule, and number of jointly trained datasets.
+We therefore leave systematic scaling (e.g., expert count/width, per-stage sharing, alternative routers) to future work.
+Finally, because 3D semantic segmentation depends on costly human labels, we plan to pursue self-supervised pretraining on diverse unlabeled point clouds to reduce annotation effort and improve generalization.
+We hope these results serve as a simple, strong baseline and a useful reference point for future works.
 A.2 Artifacts and License
 We report a list of licenses for all datasets and code used in our experiment in Tab.
-4.
+6.
 We strictly follow all the datasets and code licenses and limit the scope of these datasets and code to academic research only.
@@
 Appendix B Additional Experiments
-B.1 Experiment Settings
-Data.
+B.1 Experiment Datasets
 We compare a diverse set of 3D semantic segmentation datasets spanning both indoor and outdoor environments.
@@
 We include detailed dataset statistic in Tab.
-5.
+7.
 Furthermore, we provide detailed per-class comparisons on semantic classes for both indoor and outdoor datasets, as shown in Fig. 10 and Fig. 10.
@@
 nuScenes [9] 32-beam LiDAR L=[-105.17, 105.25], W=[-105.30, 105.17], H=[-53.42, 20.01] outdoor
-Training.
-Our default joint training strategy and Point-MoE-L configs are summarized in Tab.
-8.
-We experiment with two batching schemes: (1) single batch, where each iteration samples points exclusively from one dataset, and (2) mix batch, where samples are drawn from multiple datasets with equal probabilities.
-We find that larger batch sizes lead to more stable expert assignment and better convergence behavior in the Mixture-of-Experts architecture.
-All models are trained from scratch on the joint multi-domain datasets without any fine-tuning stage.
-Table 6: Categorical settings of Indoor Datasets.
+Table 8: Categories of Indoor Datasets.
 Dataset #C
@@
 Matterport3D 21
-Table 7: Categorical settings of Outdoor Datasets.
+Table 9: Categories of Outdoor Datasets.
 Dataset #C
@@
 domain rand. False
-iters 120k iters 160k expert capacity 2H
+iters 140k iters 180k expert capacity 2H
 (a) Detector trained on indoor domains
@@
 However, our findings with Point-MoE offer a different perspective: by explicitly modeling distributional variation through expert specialization, the architecture may naturally accommodate such domain shifts without requiring aggressive data homogenization.
-Does Point-MoE work on single dataset?
-Table 9: Performance of Point-MoE vs.
-PTv3 on individual datasets.
-We evaluate both models independently on each dataset to examine whether Point-MoE offers benefits even without cross-domain training.
-Point-MoE performs comparably to PTv3.
-Method ScanNet Structured3D Matterport3D nuScenes Waymo Average
-PTv3 75.01 79.37 53.89 77.16 70.20 71.13
-Point-MoE 75.63 79.29 53.47 78.24 69.50 71.23
-We evaluate Point-MoE without cross-domain supervision by training and testing independently on three indoor datasets (ScanNet [21], Structured3D [88], Matterport3D [12]) and two outdoor datasets (nuScenes [10], Waymo [65]).
-On ScanNet and nuScenes, Point-MoE outperforms PTv3 by +0.62 and +1.08, respectively.
-On Structured3D (evaluated on its test split), performance is nearly identical (–0.08).
-On Matterport3D (test split) and Waymo (validation split), Point-MoE underperforms PTv3 by –0.42 and –0.70, respectively.
-On average, Point-MoE achieves 71.23 vs.
-71.13 for PTv3 (+0.10).
-These results suggest that expert specialization can emerge in single-domain settings, but the overall improvement is small and not consistent across datasets.
-We consider the findings inconclusive.
-We hypothesize that Point-MoE may require different configurations or training strategies when applied to a single dataset.
-In particular, the current gating and load balancing mechanisms are optimized for multi-domain settings, where diverse token distributions benefit from expert diversity.
-In contrast, single-domain training may benefit from fewer experts, a larger top-$k$, or stronger regularization to avoid expert under-utilization.
-Table 10: Joint Training Results on Indoor 3D Semantic Segmentation without Precise Evaluator.
-Each model is trained jointly on three indoor datasets and evaluated under a domain-agnostic setting without relying on ground-truth domain labels during inference.
-We report performance when models are required to predict across all domains.
-To evaluate generalization with a single “generic“ label, results marked (G) use this label for both seen and unseen domains.
-Unmarked results rely on domain detector predictions..
-Methods Params Activated Mix Batch Layer. ScanNet Structured3D S3DIS Matterport3D
-Val Val Val Val
+Does Mixture-of-Experts (MoE) work with a sparse convolutional backbone?
+Table 11: MoE on a sparse convolutional backbone (OA-CNN).
+Numbers are mIoU.
+Rows (1)–(4) use the public OA-CNN implementation (with a language-guided head for cross-dataset evaluation).
+(1)–(3) train on a single dataset; (4) joint-trains on ScanNet, Structured3D, and S3DIS; (5) replace SparseCNN with MoE-CNN using top-2 routing.
+Model ScanNet Structured3D S3DIS Matterport3D Average
+(1) OA-CNN-S3DIS 5.70 3.8 65.3 5.4 20.6
+(2) OA-CNN-ScanNet 72.6 21.6 8.4 37.4 35.0
+(3) OA-CNN-Structured3D 20.8 65.6 6.8 18.9 28.0
+(4) OA-CNN 71.6 59.0 60.6 37.8 57.3
+(5) OA-CNN-MoE 72.0 65.3 63.0 40.2 60.1
+Yes.
+We integrate MoE into OA-CNN [52] by replacing the SparseCNN layer in the BasicBlock with a mixture of sparse-convolution experts (MoE-CNN), evaluating with top-2 routing.
+As summarized in Table 11, MoE improves over the joint-trained OA-CNN baseline on seen domains (ScanNet, Structured3D, S3DIS) and the unseen Matterport3D, raising the average from 57.3 to 60.1.
+This demonstrates that MoE is compatible with sparse convolutional backbones and delivers measurable gains, which further demonstrates MoE’s applicability to architecture beyond transformers.
+Table 12: Main Results without Precise Evaluator.
+Each model is trained jointly on three indoor datasets and evaluated without precise evaluator.
+* indicates that each training batch combines samples from multiple datasets, methods with ** use mix dataset batch and layernorm [3].
+Methods Params Activated ScanNet Structured3D S3DIS Matterport3D Average
+Val Val Area5 Val
 Single-dataset Training
-PTv3-ScanNet [78] 46M 46M 74.4 19.4 8.5 37.9
-PTv3-Structured3D [78] 46M 46M 20.0 78.0 7.0 19.5
-PTv3-S3DIS [78] 46M 46M 4.9 3.4 66.1 5.2
-PTv3-Matterport3D [78] 46M 46M 52.7 23.0 9.5 53.6
-Domain-agnostic joint training
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-Domain-aware joint training
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-Table 11: Cross-Domain 3D Semantic Segmentation on Indoor and Outdoor Scenes without Precise Evaluator.
-Models are trained jointly on both indoor and outdoor datasets under domain-agnostic supervision and domain-aware joint training settings.
-(G): Using the “generic“ label during evaluation (see definition above).
-Methods Params Activated Mix Layer. ScanNet Structured3D S3DIS nuScenes SemKITTI Matterport3D Waymo
-Val Val Val Val Val Val Val
-Domain-agnostic joint training
-[…]
-[…]
-[…]
-[…]
-Domain-aware joint training
-[…]
-[…]
-[…]
-[…]
-Table 12: Joint Training Performance on Indoor 3D Semantic Segmentation with Ground-Truth Domain Labels.
-We additionally evaluate models trained on indoor datasets using domain-aware supervision, where explicit ground-truth domain labels are provided during training to guide the learning process.
-Methods Params Activated Mix Batch ScanNet Structured3D S3DIS
-Val Val Test Val
-Domain-aware joint training with G.T. labels
-PPT-S [80, 74] 47M 47M 74.7 60.9 65.5 71.6
-PPT-L [80, 74] 98M 98M 74.6 60.9 62.6 69.2
-Point-MoE-S 59M 52M 76.7 66.7 72.3 71.9
-Point-MoE-L 100M 60M 75.6 69.7 71.3 69.6
-Point-MoE-G-S 59M 52M 75.7 69.1 70.8 68.1
-Point-MoE-G-L 100M 60M 76.7 68.1 69.6 67.1
-Table 13: Joint Training Performance on Indoor-Outdoor 3D Semantic Segmentation with Ground-Truth Domain Labels.
-We evaluate models trained on indoor-outdoor datasets using domain-aware supervision with ground truth labels.
-Methods Params Activated Mix ScanNet Structured3D S3DIS nuScenes SemKITTI
-Val Val Test Val Val Val
-Domain-aware joint training with G.T. labels
-PPT-L [80, 74] 98M 98M 76.7 67.2 66.0 71.2 72.6 66.9
-Point-MoE-L 100M 60M 76.8 69.6 70.1 70.6 72.4 64.5
-Point-MoE-G-L 100M 60M 77.0 74.9 73.2 71.6 73.0 64.3
+PTv3-ScanNet [81] 46M 46M 74.4 19.4 8.5 37.9 35.1
+PTv3-Structured3D [81] 46M 46M 20.0 78.0 7.0 19.5 31.1
+PTv3-S3DIS [81] 46M 46M 4.9 3.4 66.1 5.2 19.9
+PTv3-Matterport3D [81] 46M 46M 52.7 23.0 9.5 53.6 34.7
+Multi-dataset Joint Training
+PTv3-[81] 46M 46M 48.4 57.8 44.1 27.8 44.5
+PTv3-S* [81] [81] 46M 46M 67.7 45.9 62.5 34.9 52.8
+PTv3-S** [81] 46M 46M 70.0 56.0 66.9 35.6 57.1
+PTv3-[81] 97M 97M 57.1 58.1 40.3 31.0 46.6
+PTv3-L* [81] 97M 97M 73.3 57.5 67.2 38.4 59.1
+PTv3-L** [81] [81] 97M 97M 74.0 67.4 67.8 40.6 62.5
+PPT-[82, 75] 47M 47M 74.2 58.2 61.0 24.3 54.4
+PPT-[82, 75] 98M 98M 74.7 64.9 62.0 27.8 57.4
+Point-MoE-S 59M 52M 74.5 66.9 68.6 40.1 62.5
+Point-MoE-L 100M 60M 75.1 69.3 68.5 41.3 63.6
+Table 13: Main Results on Indoor and Outdoor Scenes without Precise Evaluator.
+Models are trained jointly on both indoor and outdoor datasets.
+indicates that each training batch combines samples from multiple datasets, methods with ∗∗ use mix dataset batch and layernorm [3].
+Methods Params Activated ScanNet Structured3D S3DIS nuScenes SemKITTI Matterport3D Waymo Average
+Val Val Area5 Val Val Val Val
+PTv3-L [81] 97M 97M 55.9 47.6 19.3 42.9 35.3 31.8 12.4 35.0
+PTv3-L* [81] 97M 97M 73.4 61.9 61.4 64.7 59.3 38.9 22.8 54.6
+PTv3-L** [81] 97M 97M 73.6 61.3 65.1 64.5 59.7 37.1 21.9 54.7
+PPT-L [82, 75] 98M 98M 75.2 67.5 61.0 57.2 64.0 20.8 16.0 51.7
+Point-MoE-L 100M 60M 76.2 70.1 67.9 68.9 63.3 40.7 23.4 58.6
 B.3 Additional Main Results
@@
 The results reported in Tab.
-1 and Tab.
 1 are obtained using the precise evaluator provided by Pointcept [18].
@@
 12 and Tab.
-2, we observe that Point-MoE outperforms the baselines in most domains in both the indoor-only and indoor-outdoor joint training settings.
+13, we observe that Point-MoE outperforms the baselines in most domains in both the indoor-only and indoor-outdoor joint training settings.
 These results are consistent with those obtained using the precise evaluator reported in the main text.
@@
 As such, our analysis and conclusions align closely with those presented in the main text.
-How Strong is Domain-Aware Point-MoE?
-In the main text, we compare domain-agnostic and domain-aware settings and observe that domain-aware Point-MoE does not consistently outperform its domain-agnostic counterpart.
-Here we want to offer discussions and analysis to explain the reason.
-One contributing factor is the reliance on the domain predictor as we detailed in sec.
-B.2.
-Although the domain predictor achieves high classification accuracy, even minor errors can lead to noticeable performance degradation—either by misrouting inputs to suboptimal experts in Point-MoE or by selecting incorrect normalization statistics in PPT.
-To better isolate this effect, we include results in Tab.
-12 and Tab.
-13, where we evaluate domain-aware Point-MoE using ground-truth domain labels.
-In general, Point-MoE continues to outperform standard baselines in both settings.
-Moreover, we find that domain-aware Point-MoE with access to ground-truth labels achieves better performance on seen domains, while domain-agnostic Point-MoE trained with domain-randomized data generalizes better to unseen domains.
-Interestingly, we also observe that Point-MoE trained with generic (i.e., domain-agnostic) supervision performs well even on seen domains when exposed to more diverse training inputs.
-This suggests that domain-randomization during training may benefit performance across both seen and unseen domains in challenging settings.
-Finally, we note that domain-aware Point-MoE can yield stronger results when accurate domain labels are available at inference time.
+To ensure a strong baseline, we add mixed-dataset training and tune hyper-parameters, reporting the baseline with its best configuration.
+To be specifically, we apply mix-dataset training for baselines we can directly apply without code modification and we also find layernorm is empirically better for the baselines.
 Appendix C Additional Visualization and Discussions on Point-MoE
-C.1 t-SNE
-Figure 6: t-SNE Visualization.
-The first three columns illustrate feature similarities and differences across PTv3, PPT, and Point-MoE for both encoder and decoder layers.
-The rightmost column highlights how Point-MoE generalizes to unseen domains, with Matterport3D features aligning closely with the most relevant seen domains—demonstrating robust cross-domain generalization.
-Fig.
-6 presents t-SNE visualizations [69] of encoder and decoder features from PTv3, PPT [80], and Point-MoE.
-PTv3 exhibits no clear separation in either encoder or decoder spaces, indicating its inability to structure domain-specific representations.
-In contrast, PPT shows strongly separated clusters for each domain at both encoder and decoder levels, suggesting it overfits to domain-specific patterns.
-While this benefits seen domains, it may hinder generalization to unseen settings.
-Point-MoE, however, reveals a more interesting behavior.
-Its encoder features remain entangled across domains—indicating shared representation learning—while its decoder learns to disentangle domain-specific structures.
-This division of labor suggests that Point-MoE effectively performs domain inference within the model, without requiring explicit domain labels.
-Most notably, in the unseen Matterport3D domain, Point-MoE maps samples to nearby clusters of the most relevant seen domains (e.g., ScanNet), enabling strong generalization.
-This implicit alignment underscores the model’s ability to organize and interpret new environments by leveraging expertise learned from diverse sources.
-C.2 Semantic Concept Distribution over Experts
+C.1 Semantic Concept Distribution over Experts
 Expert Specialization Word Cloud.
@@
 From the visualization in Fig.
-7, we observe several notable trends.
+6, we observe several notable trends.
 Specialization emerges clearly: within each decoder block, experts tend to focus on semantically coherent regions such as bed, bathtub, window, or trailer, suggesting that experts develop functional roles under shared supervision.
@@
 To complement the word cloud analysis, we provide fine-grained views of expert behavior in Fig.
+7, Fig.
 8, Fig.
-9, Fig.
 9 and Fig.
-11.
+10.
 For each semantic class, we visualize the soft routing probabilities across all experts within a specific decoder block.
@@
 Figure 10: Expert Choices Visualization for Decoder Layer 3.
-C.3 Expert Evolution.
+C.2 Expert Evolution.
 Figure 11: JSD Dynamics During Training.
@@
 In Fig.
-12, we track the Jensen-Shannon Divergence (JSD) between expert assignment distributions across different datasets every 10k training iterations to analyze how expert specialization evolves over time.
+11, we track the Jensen-Shannon Divergence (JSD) between expert assignment distributions across different datasets every 10k training iterations to analyze how expert specialization evolves over time.
 Let $P_{1},P_{2},\dots,P_{n}$ denote the expert routing distributions for each of the $n$ datasets, where each $P_{i}$ is a categorical distribution over experts.
@@
 Additional visualizations of expert assignments can be found in Sec.
-C.4, further illustrating this specialization.
-C.4 More Expert Choice Visualization
+C.3, further illustrating this specialization.
+C.3 More Expert Choice Visualization
 Fig.
+12, Fig.
 13, Fig.
@@
 16, Fig.
-17, Fig.
-18, and Fig.
+17, and Fig.
 18 showcase expert assignments for a validation scene across all layers.
@@
 In Fig.
-13, we observe that early encoder layers primarily rely on geometric cues for expert routing.
+12, we observe that early encoder layers primarily rely on geometric cues for expert routing.
 For example, certain experts—such as the green one—are consistently activated along object boundaries, including the edges of desks and chairs, while others tend to dominate on flat or horizontal surfaces.
@@
 In Fig.
-14, we observe a similar overall pattern to ScanNet [21].
+13, we observe a similar overall pattern to ScanNet [19].
 In the encoder, expert assignments are primarily driven by local geometric structure, with clear attention to furniture boundaries and wall intersections.
@@
 In Fig.
-17, we examine an outdoor scene from SemanticKITTI [6], characterized by sparse LiDAR data.
+16, we examine an outdoor scene from SemanticKITTI [5], characterized by sparse LiDAR data.
 Despite the limited geometric density, the model still exhibits meaningful expert routing: nearby points tend to be assigned to certain experts, while distant points activate others.